### v1

# Cosmological Zoom-In Simulations of Milky Way Host Size Dark Matter Halos with a Blue-Tilted Primordial Power Spectrum

Abstract
Recent observations from the James Webb Space Telescope revealed a surprisingly large number of galaxies formed at high redshift. Along with strong lensing studies and nearby galaxy observations, these could challenge the standard Lambda Cold Dark Matter cosmology with a power-law primordial power spectrum. In this study, we conduct high-resolution cosmological zoom-in dark matter-only simulations of Milky Way host size halos with a blue, tilted primordial power spectrum ($P(k)\propto k^{m_{s}}$ with $m_{s}>1$ at small scales $>1~{}{\rm Mpc}^{-1}$). We find that the blue-tilted subhalo mass functions can be enhanced by more than a factor of two for subhalo masses $M_{\rm sub}\lesssim 10^{10}~{}{\rm M}_{\odot}$, whereas the subhalo $V_{\rm max}$ functions can be enhanced by a factor of four for maximum circular velocities $V_{\rm max}\lesssim 30~{}{\rm km/s}$. The blue-tilted scaled cumulative substructure fraction can be an order of magnitude higher at $\sim$10of the virial radius. The blue-tilted subhalos also have higher central densities, since the blue-tilted subhalos reach the same $V_{\rm max}$ at a smaller distance $R_{\rm max}$ from the center. We have also verified these findings with higher-resolution simulations.
I Introduction
The standard cosmology model includes the single-field slow-roll inflation model Guth (1981); Liddle and Lyth (1993); Steinhardt and Turner (1984); Salopek and Bond (1990); Liddle and Lyth (1992), which predicts the approximately single power-law (PL) primordial power spectrum (PPS) (Lyth and Riotto, 1999), and the lambda cold dark matter model (LCDM, Peebles (1993); Frenk et al. (1985)), which governs the later evolution of our universe. It has achieved great successes on large scales, e.g. supported by the cosmic microwave background (CMB) Planck Collaboration et al. (2020a, b), galaxy surveys Troxel et al. (2018); Blanton et al. (2017) and Lyman-alpha forest Chabanier et al. (2019). However, there are relatively sparse constraints for the scales smaller than $\sim 1~{}{\rm Mpc}^{-1}$. In particular, there have been active debates over the so-called small-scale crisis of LCDM Bullock and Boylan-Kolchin (2017), including the Missing Satellite Problem (MSP, (Klypin et al., 1999; Moore et al., 1999)), the Core-Cuspy problem (Salucci and Burkert, 2000; Flores and Primack, 1994), and the Too-Big-To-Fail problem (Boylan-Kolchin et al., 2011), etc. This leaves room for other models beyond the standard cosmology, including different models of dark matter, effects of baryons, and modifications to the PPS.
One possible modification is to enhance the PPS on small scales (${\rm>1~{}{\rm Mpc}^{-1}}$) (Covi and Lyth, 1999; Martin and Brandenberger, 2001; Gong and Sasaki, 2011), deviating from the PL PPS. In these small-scale enhanced models, PPS will follow the traditional PL model on large scales, preserving the success of the PL PPS, while having an enhancement on small scales. For example, Ref. Hirano et al. (2015) considered blue-tilted (BT) models, which are broken power law PPSs with a higher power law index in scales $>1~{}{\rm Mpc}^{-1}$. They found a significant boost in dark matter halos with BT BBS, especially at high redshift. More complicated forms also exist, such as several bumps in the PPS Tkachev et al. (2024).
With the recently launched James Webb Space Telescope (JWST) Gardner et al. (2006), Ref. Labbé et al. (2023) found a large number of galaxies formed in the high redshift universe ($z>10$), which may be in tension with the standard cosmology Boylan-Kolchin (2023); Lovell et al. (2023). Ref. Parashari and Laha (2023) eased this tension with the BT formalism, based on the Press-Schechter formalism Press and Schechter (1974) and the Sheth-Tormen correction Sheth and Tormen (1999). Ref. Hirano and Yoshida (2024) performed cosmological N-body simulations with the BT model, achieving similar conclusions that the BT model could match the JWST observations without invoking ultra-high star formation rates. However, other mechanisms in galaxy formation models (e.g. Refs. Dekel et al. (2023); Haslbauer et al. (2022); Yung et al. (2024)) could potentially explain the JWST data. Thus, apart from the high redshift massive galaxies, additional observations are required to distinguish the BT and PL models.
The strong lensing observations can provide another constraint. The asymptotic flux ratio relation, which serves as a general principle in smooth lensing potentials (Mao, 1992), is often observed to be violated. This is taken as evidence for the presence of the substructures around lensing galaxies Mao and Schneider (1998). However, in the PL LCDM cosmology, there could be not enough low mass substructures in lensing galaxies to explain the violations Macciò and Miranda (2006); Xu et al. (2009, 2015). This could suggest extra structures not present in dark matter-only simulations under the standard PL LCDM cosmology. On the other hand, Ref. Gilman et al. (2022) constrained sub-galactic PPS to be consistent with PL with the 11 strongly lensed and quadruply image quasars, albeit with possibilities with blue/red tilted PPS on small scales.
The density profiles and statistics of nearby low-mass dwarf galaxies can also inform us about the nature of dark matter, baryonic physics, and the PPS. For example, the observed central densities of dwarf galaxies can provide insights into the nature of dark matter Rocha et al. (2013); Yang et al. (2024) or constraining baryonic physics Di Cintio et al. (2014); Chan et al. (2015); Tollet et al. (2016); Lazar et al. (2020). The “missing satellite problem” describes the discrepancy between the number of subhalos from cosmology dark matter simulations and the observed satellite galaxy population (Klypin et al., 1999; Moore et al., 1999). This problem could signify a challenge to the standard PL LCDM cosmology, or the importance of including accurate galaxy formation physics in dwarf galaxies. Nearby dwarf galaxies also provide constraints on the PPS. For example, Refs. Esteban et al. (2024); Dekker and Kravtsov (2024) employed semi-analytic models to constrain the BT PPS parameter space, using the observed relation between half-light radius and V-band luminosity function in nearby dwarf galaxies.
Thanks to the recent improvements in observations, many more ($\gtrsim 50$) faint satellite galaxies have been discovered in Milky Way (MW) and nearby MW mass hosts Drlica-Wagner et al. (2015); Bechtol et al. (2015); Koposov et al. (2015); Homma et al. (2024); Müller et al. (2024). After correcting for the detection efficiency (completeness correction), Ref. Kim et al. (2018) argued that the “missing satellites problem” can be solved. It could even become the “too many satellites problem”, meaning that there are too many satellite galaxies inferred from observations than expected in simulations/models Kim et al. (2018). This happens if some mechanisms can reduce the number of satellites in simulations /models. For example, a baryonic disk can tidally strip subhalos Garrison-Kimmel et al. (2017), and/or the minimum subhalo mass to host an observable galaxy is large Graus et al. (2019). This “too many satellites problem” motivates us to explore mechanisms to boost the number of subhalos, which includes the BT model.
Motivated by high redshift, strong lensing, and nearby galaxy observations, we explore the impact of the BT PPS on the substructures in dark matter halos, which, to our knowledge, has not been previously investigated with numerical simulations. We perform cosmological zoom-in dark matter-only simulations of MW host size halos with BT PPS to study its substructures at $z=0$. Note that we only modify the PPS but still consider only cold dark matter (i.e. not warm or self-interacting dark matter).
In section II, we will give a theoretical overview, laying down the framework for the BT PPS and our simulation. Then in section III, we introduce our numerical codes for initial conditions, cosmological simulations, and halo identification. We present our main results in section IV and conclude in section V. Additionally, as the first work to use HBT-HERONS (see subsection III.3) for zoom-in simulations, we show the comparison between HBT+HERONS and another halo finder VELOCIraptor in Appendix A. We also run higher-resolution simulations and compare the fiducial results in Appendix B.
II Theoretical Overview
Figure 1: The conceptual steps of the cosmological stages and our simulation steps. The upper panel shows the universe’s chronology, along with redshifts of some critical eras; the lower panel shows the numerical tools we used.
The upper panel of Figure 1 shows the cosmic chronology. Shortly after the Big Bang, the primordial power spectrum (PPS) and small inhomogeneities were established during the inflation era. At the cosmic recombination $z\sim 1100$, electrons are combined with protons to form neutral hydrogens, allowing radiation to travel freely and resulting in the cosmic microwave background (CMB). Under the influence of gravity, the small density fluctuations are amplified, which can still be described by Lagrangian perturbation theories. By $z\sim 100$, non-linearities started to grow and more dark matter collapsed into halos, requiring N-body simulations to model accurately. Eventually, stars and galaxies are born at the centers of dark matter halos, leading to the structure we observe today. Our study connects the very early universe to the present-day nearby galaxy structures and statistics.
The matter power spectrum in the post-recombination era is connected to the primordial power spectrum $P_{i}(k)$ through Mo et al. (2010):
$P(k,t)=P_{i}(k)T^{2}(k)D^{2}(t),$ (1)
where $T(k)$ is the linear transfer function and $D(t)$ is the growth factor. In the traditional single-field slow-roll inflation, the PPS follows the PL model:
$P_{i}(k)\propto k^{n_{s}},$ (2)
with the spectral index $n_{s}\sim 0.96$ (see subsubsection III.2.1).
Ref. Hirano et al. (2015) gave the following formalism for the BT models:
$P_{i}(k)\propto\left\{\begin{matrix}k^{n_{s}},&({\rm for}\;k\leq k_{p}),\\ k^{n_{s}}\cdot{(\frac{k}{k_{p}})}^{m_{s}-n_{s}},&({\rm for}\;k>k_{p}),\\ \end{matrix}\right.$ (3)
which is a broken power law modification of Equation 2. We introduce $k_{p}$ as the pivot scale and $m_{s}$ as the spectral index on small scales. We require $m_{s}$ to be larger than $n_{s}$ to give an enhancement (i.e. blue-tilted, BT). When $m_{s}$ is smaller than $n_{s}$, the corresponding PPS model is called the red tilted (RT) model Dekker and Kravtsov (2024).
In this paper, along with the PL model, we pick two BT models from Ref. Hirano and Yoshida (2024) [^1] For $k_{p}$, the value in our table is a float number to several digits because original models use $h~{}\rm{Mpc}^{-1}$ as a unit of $k_{p}$, but we convert it into ${\rm Mpc}^{-1}$.: BT_deep and BT_soft, listed in Table 1. BT_soft with a reasonable star formation efficiency $\sim 0.2$ can explain the large number of galaxies formed at high redshift as observed by JWST Hirano and Yoshida (2024). However, to match the JWST observations, BT_deep requires a high star formation efficiency, comparable to PL Hirano and Yoshida (2024). Although BT_deep does not help to explain the JWST observations, it is still interesting to examine its impact on the substructures of dark matter halo.
Both models are within the allowable parameter space in Ref. Dekker and Kravtsov (2024) (constrained with nearby dwarf central densities), although BT_soft is the marginal case. We also confirm that our models are within the parameter space constrained with strong lensing Gilman et al. (2022).
For all of our PPS models, the matter power spectra at z=1089 and the parameter settings are shown in Figure 2 and Table 1 respectively. BT_soft deviates from PL at $0.7~{}{\rm Mpc}^{-1}$, whereas BT_deep deviates at $3.51~{}{\rm Mpc}^{-1}$, but they have the same power law index afterward.
Figure 2: The upper panel shows the power spectra for the matter density perturbation at z=1089 with PL (orange), BT_deep (green), and BT_soft (red). The inset is for a wider wave vector $k$ range. The bottom panel shows the ratios of the BT to PL power spectrum. The model parameters are listed in Table 1.
Models Related parameters
PL Power Law Primordial Power Spectrum
$n_{s}=0.961$
BT_deep $k_{p}=3.51\ \rm{Mpc}^{-1}$ $m_{s}=1.5$
BT_soft $k_{p}=0.702\ \rm{Mpc}^{-1}$ $m_{s}=1.5$
Table 1: The parameters of all the chosen models. $k_{p}$ is the wave vector at which the BT PPS would deviate from the PL PPS. $m_{s}$ is the enhanced spectral index for $k>k_{p}$, at the small scales. For other cosmological parameters, see subsubsection III.2.1.
The wave number $k$ can be connected with a mass scale $M_{l}$ at z=0 Bullock and Boylan-Kolchin (2017), by assuming the sphere radius $r_{l}$ to be half wavelength $\frac{\pi}{k}$:
$M_{l}=$ $\frac{4\pi}{3}{r_{l}}^{3}\rho_{m}=\frac{\Omega_{m}H_{0}^{2}}{2G}r_{l}^{3}$
$=$ $1.71\times 10^{11}\left(\frac{\Omega_{m}}{0.3}\right)\left(\frac{H_{0}}{70}\right)^{2}\left(\frac{r_{l}}{1~{}\rm{Mpc}}\right)^{3}~{}{\rm M}_{\odot}.$ (4)
We further give the pivot mass $M_{p}$ under which the BT model can enhance the structure formation:
$M_{p}=5.29\times 10^{12}\left(\frac{\Omega_{m}}{0.3}\right)\left(\frac{H_{0}}{70}\right)^{2}\left(\frac{k_{p}}{1~{}{\rm Mpc}^{-1}}\right)^{-3}~{}{\rm M}_{\odot}.$ (5)
For the BT_deep and BT_soft models, $M_{p}$ would be $\sim 1.1\times 10^{11}~{}{\rm M}_{\odot}$ and $\sim 1.4\times 10^{13}~{}{\rm M}_{\odot}$ respectively, using the cosmological parameters in our paper (see subsubsection III.2.1). The host halo of an MW-size galaxy is believed to be around $M_{\rm halo}\sim 10^{12}~{}{\rm M}_{\odot}$. Therefore, while both the BT_deep and BT_soft models can enhance the substructures of MW host size halos, only in BT_soft the main halo can be affected.
To illustrate the impact of the BT model on the dark matter halo statistics across different redshifts, we apply a public semi-numerical code genmf Reed et al. (2007)  [^2] Publicly available at https://icc.dur.ac.uk/Research/PublicDownloads/genmf_readme.html. It can predict halo mass functions in cosmological simulations at a given redshift. Ref. Hirano and Yoshida (2024) verified that genmf is accurate within a factor of 2-3, by comparing it with numerical N-body simulations (with PL and BT).
We use genmf to show the cumulative halo mass function at $z=0$ and $z=9$ for the BT and PL models in Figure 3. The figure shows, at $z=9$, the BT models can produce an order-of-magnitude more halos, especially for $M_{\rm halo}<10^{10-11}{\rm M}_{\odot}$. At $z=0$, the enhancement in the number of halos reduces to a factor of 2-3 for $M_{\rm halo}<10^{10-11}{\rm M}_{\odot}$. Ref. Reed et al. (2007) also found that the high redshift halo mass function is more sensitive to cosmology, due to the stiffness of the function. This high redshift enhancement motivates Ref. Hirano and Yoshida (2024) to explain the early formation of galaxies discovered by JWST with the BT models.
However, at both redshifts, there is no enhancement at the high-mass end, since the BT models can only enhance the structures lower than the pivot mass $M_{p}$ (Equation 5), shown as the triangles in the figure.
While genmf can estimate the halo mass functions in the BT models within a factor of two, we still require cosmological N-body simulations to accurately capture the sub-halo distributions and statistics in the BT models. This will help us explore BT PPS’s impact on nearby galaxies and strong lensing observations.
Figure 3: The cumulative halo mass function generated by genmf, at z=9 (dashed line) and z=0 (solid line), with PL (orange), BT_deep (green), and BT_soft (red). The upper panel shows the cumulative halo mass function per unit comoving volume. The bottom panel shows the ratios of BT to PL. The triangle symbols on the solid lines show the pivot masses $M_{p}$ for BT_deep (green) and BT_soft (red), respectively.
III Numerical Methods
Our cosmological simulation pipeline is outlined in the bottom panel of Figure 1. We will introduce more details in the following three subsections.
III.1 Zoom-in initial conditions (IC)
We carry out a suite of cosmological zoom-in simulations of MW host-size halos. We increase the number of particles in a selected zoom-in region, which contains high-resolution (HRS) particles. The particles outside the zoom-in region are called background or low-resolution (LRS) particles. We chose the public code MUSIC - multi-scale cosmological initial conditions (Hahn and Abel, 2011), which can handle multi-scale IC, to generate initial conditions for our cosmological zoom-in simulations. We adopted the Eisenstein and Hu (Eisenstein and Hu, 1998) transfer function and modified it according to section II to mimic the blue-tilted primordial power spectra.
For the convenience of comparison, we adopt the MUSIC configuration files of the existing simulation project FIRE (Hopkins et al., 2018; Wetzel et al., 2023). We choose the zoom-in initial conditions at the size of MW host halos (i.e., $M_{h}\sim 10^{12}~{}{\rm M}_{\odot}$ in mass), specifically, m12i (Wetzel et al., 2016) and m12f (Garrison-Kimmel et al., 2017). We obtain those configuration files from the Flatiron Institute Data Exploration and Comparison Hub  [^3] Publicly available at https://flathub.flatironinstitute.org/fire.
III.2 Simulation code
We perform several zoom-in simulations using the state-of-the-art cosmological simulation code SWIFT (Schaller et al., 2018, 2024) [^4] Publicly available at https://swiftsim.com/. For N-body simulations, SWIFT uses the Fast Multipole Method (FMM) at small scales (Dehnen, 2000), which can reduce the complexity of gravity calculation (including both the potential and force calculation) to $O(N)$. The Particle Mesh Method is also coupled at large scales to handle periodic volumes (Hockney and Eastwood, 1988).
III.2.1 Cosmological parameters
For all the cosmological simulations in this paper, we adopt the WMAP-7 result (Komatsu et al., 2011), which is also used in the original m12i (Wetzel et al., 2016) and m12f (Garrison-Kimmel et al., 2017) simulations.
  1. 1.
    Universe density parameters $\Omega_{m}=0.272$, $\Omega_{b}=0.0455$, $\Omega_{L}=0.728$.
  2. 2.
    Hubble constant $H_{0}=70.2\ \rm{km}\ \rm{s}^{-1}\ \rm{Mpc}^{-1}$
  3. 3.
    Scalar spectral index $n_{s}=0.961$
  4. 4.
    Root-mean-square matter fluctuation averaged over a sphere of radius $8h^{-1}~{}\rm{Mpc}$, $\sigma_{8}=0.807$.
III.3 Halo finder and halo property post-processor
Current halo finders can be categorized into three types based on working mechanism: (A) the configuration space finders, (B) the phase space finders (which utilize both spatial and velocity space information), and (C) the tracking finders (which build the particle list of a certain subhalo based on its progenitors).
In the main text of the paper, we adopt HBT+HERONS (Han et al., 2017) (Forouhar Moreno et al in preparation)  [^5] This is an updated version of HBT+ that improves the tracking of subhalos. It can be found at https://github.com/SWIFTSIM/HBT-HERONS as our halo finder, which is a tracking finder. We also explore the result with a phase space finder, VELOCIraptor (Elahi et al., 2019), and compare their results in Appendix A. In this paper, we adopt 20 as the minimum number of particles to be identified as a dark matter halo for both halo finders. We choose the position of the most bound particle as the halo center.
We use SOAP (Spherical Overdensity and Aperture Processor) [^6] https://github.com/SWIFTSIM/SOAP to calculate the spherical overdensity properties, based on the results from HBT+HERONS.
III.3.1 Convergence radius
The dark matter halo’s radial density profile would be underestimated within a resolution limit radius in N-body simulations, due to the two-body relaxation effect Power et al. (2003). This would artificially turn a cuspy profile into a core profile. This radius is called the convergence radius and defined as the smallest $r$ fulfilling Equation 6:
$\frac{t_{\rm relax}(r)}{t_{0}}=\frac{\sqrt{200}}{8}\frac{N(r)}{\ln N(r)}\left[\frac{\rho(r)}{\rho_{\rm crit}}\right]^{-1/2}>0.6,$ (6)
where $\rho_{\rm crit}$ is the critical density of universe, $\rho(r)$ is the average density within radius $r$, $N(r)$ is the number of enclosed particles within radius $r$. $t_{0}$ is the integration time, which is close to the age of the Universe. We will indicate the regions within the convergence radius with shades in Figure 5.
IV Result
We run nine simulations using the numerical pipeline mentioned in section III. Their numerical settings are listed in Table 2.
The six fiducial-resolution simulations are used to compare the BT and the PL models. Two initial conditions, m12i and m12f, are used to validate the conclusions. Thus, it would give six fiducial-resolution simulations in total. Additionally, we run three high-resolution simulations with m12i for the resolution studies in Appendix B.
IV.1 Terminology
We give a summary of the concepts and terminologies used in our paper:
  1. 1.
    Dark Matter Halo and Subhalo: a dark matter halo is a high-density region in the universe composed of dark matter particles that are gravitationally bound to each other. Smaller overdense regions can also exist inside the halo, which are also self-bound. Those smaller halos are called the subhalos of the host halo. In zoom-in simulation, the most massive halo is called main halo, which hosts the largest number of subhalos in the zoom-in region.
  2. 2.
    $M_{\rm sub}$: the total mass of all the bound particles in a subhalo (as defined by the halo finders). In this paper, we use this mass definition for the subhalo mass function.
  3. 3.
    Spherical overdensity (SO) properties: these properties are calculated based on all the enclosed particles within a sphere around the center of the selected dark matter halo. This sphere’s radius is usually where the average density for the enclosed region reaches an integer (like 50, 100, 200, …) multiple of the critical density (labeled with 50c, 100c, 200c, …) or the background mean density (labeled with 50m, 100m, 200m, …). The properties include the spherical overdensity radius ($R_{200c},R_{200m},...$), the spherical overdensity mass ($M_{200c},M_{200m},...$), the spherical overdensity concentration parameter of the NFW fitting ($C^{NFW}_{200c},C^{NFW}_{200m},...$). With the SO radius and SO mass properties, we can calculate the SO circular velocity properties($V_{200c},V_{200m},...$) via Equation 7.
  4. 4.
    $R_{\rm vir}~{}\&~{}M_{\rm vir}$: Special kinds of the SO properties. $R_{\rm vir}$ is the virial radius of a dark matter halo. The most commonly used virial radius definition is given by Ref. Bryan and Norman (1998). $M_{\rm vir}$ is then the total mass within a sphere of radius $R_{\rm vir}$. We use this definition for the subhalo mass in subsection A.2.
  5. 5.
    $V_{\rm max},~{}V_{\rm peak}~{}\&~{}R_{\rm max}$: The circular velocity is defined as:
    $V_{\rm cir}(R)=\sqrt{\frac{GM(<R)}{R}},$ (7)
    where $R$ is the distance to the halo center, $M(<R)$ the enclosed mass that is bound to the halo, within radius $R$. The $R-V_{\rm cir}(R)$ graph is called the galaxy rotation curve, and $V_{\rm max}$ is simply the maximum circular velocity of the rotation curve. In the simulation, the subhalo would lose its mass due to the tidal forces from the host halo. However, $V_{\rm max}$ is robust and representative of a dark matter subhalo because it reflects the subhalo’s central, gravitationally bound part, which is less affected by the stripping effect and how the boundary of the subhalo is defined.
    During the evolution history of a dark matter halo, its $V_{\rm max}$ may change, and the largest $V_{\rm max}$ is called the peak circular velocity, $V_{\rm peak}$.
    The distance $R$ corresponding to the maximum circular velocity $V_{\rm max}$ is denoted as $R_{\rm max}$. For a central galaxy whose rotation curve is usually flat, $R_{\rm max}$ is quite sensitive to the simulation details even though $V_{\rm max}$ changes only slightly.
IV.2 Comparison between the PL and the BT models
Figure 4: The comparison between the dark matter mass projection maps with the PL and BT (deep/soft) PPS for m12i at z=0. The color scales are the same across the three panels. All the images are 2D plane projections of cubic volumes with a side length of $400~{}\rm{kpc}$, centered at the main halo’s most bound particle identified by HBT+HERONS.
Table 2 shows some selected properties for the main halos in all the simulations. We learn that the main halo of BT_soft is larger than that of the PL model in mass, radius, concentration, and $V_{\rm max}$. However, the BT_deep model behaves similarly to the PL model, consistent with the pivot mass $M_{p}$ based on Equation 5.
Figure 4 shows the dark matter projection maps of the PL, BT_deep and BT_soft m12i simulations. Since their sizes are similar, the main halos in the projection map look alike in these three cases. However, there are significantly more subhalos in BT than in PL, including both the small subhalos and the massive ones. For the spatial distribution, the substructures are also more centrally concentrated in BT than in PL.
Property m12i-fiducial resolution m12i-high resolution m12f-fiducial resolution
PL BT:deep BT:soft PL BT:deep BT:soft PL BT:deep BT:soft
$M_{200c}\hfill[10^{12}~{}{\rm M}_{\odot}]$ $0.950$ $0.938$ $1.12$ $0.955$ $0.945$ $1.13$ $1.32$ $1.35$ $1.50$
$M_{\rm vir}\hfill[10^{12}~{}{\rm M}_{\odot}]$ $1.14$ $1.13$ $1.28$ $1.14$ $1.11$ $1.28$ $1.61$ $1.61$ $1.73$
$M_{200m}\hfill[10^{12}~{}{\rm M}_{\odot}]$ $1.26$ $1.26$ $1.41$ $1.26$ $1.25$ $1.42$ $1.82$ $1.81$ $1.87$
$R_{200c}\hfill{\rm[kpc]}$ 202 202 214 203 202 214 226 228 236
$R_{\rm vir}\hfill{\rm[kpc]}$ 273 273 284 273 271 284 307 307 314
$R_{200m}\hfill{\rm[kpc]}$ 344 343 357 344 343 357 388 387 391
$V_{200c}\hfill{\rm[km/s]}$ 142 141 150 142 142 151 159 160 165
$C^{NFW}_{200m}$ 14.4 14.0 20.7 14.7 14.6 19.9 11.5 13.1 22.8
$V_{\rm max}\hfill{\rm[km/s]}$ 155 154 174 166 167 188 176 171 194
$m_{p}^{HRS}\hfill{\rm[10^{3}{\rm M}_{\odot}]}$ $42.3$ $42.3$ $42.3$ $5.28$ $5.28$ $5.28$ $42.3$ $42.3$ $42.3$
$l_{\rm soft}\hfill{\rm[kpc]}$ 0.208 0.208 0.208 0.104 0.104 0.104 0.208 0.208 0.208
Table 2: Some selected properties for the main halos and the basic parameters of the simulations. The first and second rows give the simulation names, encoding the main halo (m12i or m12f), the resolution levels (fiducial or high resolution), and the PPS models (PL, BT_deep or BT_soft). The rest are the properties and basic parameters in the simulations. The properties (from 3rd to 11th rows) are all explained in subsection IV.1. $m_{p}^{HRS}$ is the mass of high-resolution particle in the simulation. $l_{\rm soft}$ is the gravitational softening length (in terms of co-moving value), which is 0.02 times the mean inter-particle distance Zhang et al. (2019).
In the rest of this subsection, we would quantify these differences in terms of the subhalo mass function, the subhalo $V_{\rm max}$ function (velocity refers to $V_{\rm max}$ of the subhalos), the subhalo radial distribution (subhalo distance function), the cumulative substructure mass fraction, the subhalo $R_{\rm max}-V_{\rm max}$ relationship, and the main halo radial density profile.
IV.2.1 Main Halo Radial Density Profile
Figure 5: The density profiles of the main halos for m12i (left) and m12f (right) respectively. For both figures:
The upper panel shows the density profiles of the main halo with the PL (orange), BT_deep (green), and BT_soft (red) models; the NFW fitting line is also shown for each model as a dashed line. The bottom panel shows the ratios of the BT density to the PL density. The shaded area shows the radial range expected to be affected by the numerical two-body relaxation, given by Equation 6.
We present the radial density profiles of the main halos in Figure 5, for m12i and m12f, respectively. It is an essential property of the dark matter halo inner structure, which is well-characterized by the Navarro–Frenk–White (NFW) profile Navarro et al. (1997). We show the NFW fitting result in dashed lines and our simulation results in solid lines.
The upper panels of Figure 5 show a good match between the simulation results and the NFW fitting lines. Thus we can use the concentration parameters given by the NFW profile to describe the radial density profile. The concentration parameters of BT_deep and PL are pretty close (see Table 2), consistent with their similar radial density profiles for the main halos. Meanwhile, the BT_soft model gets a larger concentration parameter than others, consistent with its higher radial density profiles. The unresolved regions are based on the criterion described by subsubsection III.3.1.
In the bottom panels, the BT_deep model behaves like the PL model in the resolved region, while the BT_soft model is enhanced by a factor of two in the inner part of the main halo ($r<0.01R_{200m}$). It is also consistent with our calculation of the pivot masses for the deep and soft models. As the radius increases, BT and PL approach a similar density, indicating that the enhancement in the main halo density mainly occurs in the inner region.
IV.2.2 Subhalo Mass Function
We present the subhalo mass function in Figure 6. We show the subhalo count within $R_{50c}$ from the main halo center, $N(>\mu)$, as a function of $\mu={M_{\rm sub}}/{M_{200c}}$, namely the ratio between the gravitationally bound subhalo masses and the $M_{200c}$ of the main halos. We adopt this scaled mass definition because it can make the subhalo mass function insensitive to the host halo mass (Kravtsov et al., 2004; Hellwing et al., 2016). The grey “COCO fit” line is the best fitting power function for the subhalo mass ($N(>\mu)\sim\mu^{-s}$ with $s=0.95$) in Ref. (Hellwing et al., 2016), for all the subhalos within $R_{50c}$ [^7] In COCO paper, they used the empirical relation to calculate $R_{50c}$, $\sim 1.66\cdot R_{200c}$. of the main halo.
The upper panel of Figure 6 shows a reasonable match between the COCO fit and our PL simulation, showing that we are consistent with COCO. The two BT models’ lines diverge clearly from the PL line, with a higher number. The BT_deep and BT_soft subhalo mass functions almost overlap, indicating that the enhancement in subhalo mass is similar between these two models.
The flattening of the subhalo mass functions in Figure 6 is due to the limited mass resolution: we cannot resolve subhalos with fewer than $\sim 20$ particles. The shaded area shows the mass range cannot be covered due to the halo finder’s setting for the minimum halo size, which is 20 particles. Though both m12i and m12f are using the same parameters in the halo finder and thus the same mass of the minimum size halo, m12f has a larger $M_{200c}$ so its scaled subhalo mass limit $\mu=M_{\rm sub}/M_{200c}$ would be smaller.
In the bottom panel, the BT models have an apparent enhancement in the number of dark matter subhalos at different masses. Especially at the low mass range ($\mu<10^{-4},M_{\rm sub}<10^{8}~{}{\rm M}_{\odot}$), the BT effect can double the subhalo number. At the high mass range ($10^{-4}<\mu<10^{-3},10^{8}~{}{\rm M}_{\odot}<M_{\rm sub}<10^{9}~{}{\rm M}_{\odot}$), the BT models still have tens of percent enhancement. At an even larger mass ($\mu>10^{-3},M_{\rm sub}>10^{9}~{}{\rm M}_{\odot}$), the ratio curve is dominated by the noise due to a lack of statistics. We can use an inverse S shape function to fit the ratio of BT to PL, by introducing two free parameters $m,\mu_{c}$:
$f(x)=\frac{1}{2}\cdot(m-1)\cdot(\frac{-x}{\sqrt{1+x^{2}}}+1)+1,$ (8)
where
$x=4\cdot\log_{10}\frac{\mu}{\mu_{c}}.$ (9)
Here, $m$ indicates the maximum relative ratio the BT model can touch compared to the PL model, and $\mu_{c}$ is the middle point between $\mu$ starting to drop and $\mu$ value finishing the dropping. According to the definition of Equation 8, $\mu_{c}$ is also where the inverse S shape function drops to half. Since the deep and soft models perform pretty closely in the subhalo mass function, we can use one set of parameters to fit: $m=2,\mu_{c}=10^{-3}$. After getting this fitting function from the bottom ratio panel, we draw it on the upper panel of Figure 6, showing a good fit with the simulation results.
Figure 6: The cumulative scaled subhalo mass functions for m12i (colored solid) and m12f (colored dashed) respectively.
The upper panel shows the subhalo mass function for all the subhalos within $R_{50c}$ from the main halo center in the PL (orange), BT_deep (green), and BT_soft (red) simulations, with the subhalo mass $M_{\rm sub}$ scaled by $M_{200c}$ of the main halo. The bottom panel shows the ratios of the BT to PL numbers. The shaded areas show the subhalo mass cutoff due to the halo finder’s setting for minimum halo size, where we set it to 20 particles. However, based on our resolution study in Appendix B, the subhalo mass function is reliable on $\mu\geq 10^{-5.5}$, i.e. for subhalos with at least 50 particles. $\mu\sim 10^{-5.5}$ is also approximately the $\mu$ point where the subhalo mass function starts to deviate from a power-law function, i.e. a straight line in the log-log graph.
IV.2.3 Subhalo $V_{\rm max}$ Function
In Figure 7, we present the subhalo $V_{\rm max}$ function. We show the subhalo count within $R_{100c}$ from the main halo center, $N(>\nu)$, as a function of $\nu=V_{\rm max}/V_{200c}$, which is the subhalo maximum circular velocity ($V_{\rm max}$) in the units of its main halo circular velocity at $R_{200c}$ ($V_{200c}$). We adopt this scaled $V_{\rm max}$ definition because it can make the function insensitive to the main halo masses (Moore et al., 1999; Hellwing et al., 2016). The grey “Cautun fit” line results from Ref. Cautun et al. (2014) under the PL LCDM cosmology, which includes all the subhalos within $R_{100c}$ of the main halo. It is also adopted by Ref. Hellwing et al. (2016) as a fiducial reference line.
Note that when the number of particles is too few, the $V_{\rm max}$ of this subhalo is not resolved. The usual choice is not to trust subhalos with fewer than 100 particles Rodríguez-Puebla et al. (2016), which corresponds to the shaded area in the figure.
The upper panel of Figure 7 shows a good match between the Cautun fit and our PL simulation, verifying the correctness of our pipeline. While the BT_soft and BT_deep lines are higher than the PL line, the soft model has an even stronger enhancement than the deep model. It is because the BT_soft subhalos are more concentrated than the BT_deep subhalos (see subsubsection IV.2.6). Thus, the subhalo $V_{\rm max}$ function can help us to differentiate the BT_soft and BT_deep models, although the subhalo mass function cannot.
[…]
$x=4\cdot\log_{10}\frac{\nu}{\nu_{c}}.$ (10)
And we have different parameter sets for the two BT models: for BT_soft, $m=5~{}\nu_{c}=0.2$; for BT_deep, $m=3.5~{}\nu_{c}=0.15$. This indicates that the BT_soft model has a stronger enhancement for a wider range of halo sizes than the BT_deep model.
Figure 7: The cumulative scaled subhalo $V_{\rm max}$ functions for m12i (colored solid) and m12f (colored dashed) respectively. The upper panel shows the subhalo $V_{\rm max}$ function for all the subhalos within $R_{100c}$ from the main halo center in the PL (orange), BT_deep (green), and BT_soft (red) simulations, with $V_{\rm max}$ of the subhalo scaled by $V_{200c}$ of the main halo. The bottom panel shows the ratios of the BT to PL numbers. The shaded areas show the unresolved $V_{\rm max}$ range, the limit of which is the median $V_{\rm max}$ value among all the subhalos with 100 particles in the PL simulations. 100 is an empirical lower limit of particle number for dark matter halo to be resolved (Rodríguez-Puebla et al., 2016). Our resolution limit study in Appendix B also shows that for subhalos with at least 100 particles, the relative deviation in the subhalo $V_{\rm max}$ function does not exceed 30%, which is within acceptable limits.
IV.2.4 Subhalo Radial Distribution
In Figure 8, we present the subhalo radial number density profile: the normalized radial number density, $n(r)/\left<n\right>$, as a function of ${r}/{R_{200c}}$, which is the subhalo’s distance to the main halo center in the units of its main halo’s $R_{200c}$. Here, $n(r)$ is the subhalo number density in the specific radial bin; $n\right>$ is defined by:
$n\right>=\frac{N_{50c}}{\frac{4}{3}\pi R_{50c}^{3}},$ (11)
where $N_{50c}$ is the number of subhalos within $R_{50c}$ in the corressponding mass bin. The grey “Aquarius A1” line is the highest resolution simulation result of the Aquarius project Springel et al. (2008), with which the COCO project has shown its fitting. Although the original line uses kpc as a unit, we divide r over Aquarius A1’s $R_{200c}\sim 245~{}{\rm kpc}$ to get the scaled radius. Ref. Springel et al. (2008) suggested that if splitting subhalos with different mass bins across from $10^{5}~{}{\rm M}_{\odot}$ to $10^{10}~{}{\rm M}_{\odot}$, the radial distribution at every bin would all follow this line. Due to the resolution limit of our simulation, we only show the four mass bins from $10^{6}~{}{\rm M}_{\odot}$ to $10^{10}~{}{\rm M}_{\odot}$.
The upper panels of Figure 8 show the PL results between $10^{6}~{}{\rm M}_{\odot}$ and $10^{9}~{}{\rm M}_{\odot}$ fit well with the Aquarius lines. However, for the $10^{9}~{}{\rm M}_{\odot}$ to $10^{10}~{}{\rm M}_{\odot}$ mass bin, we have too few subhalos (only $\sim$ 20 within $R_{50c}$), so the plotting has a substantial scatter around the Aquarius A1 line.
In the bottom panels, for the whole mass range (or at least $10^{6}~{}{\rm M}_{\odot}\sim 10^{9}~{}{\rm M}_{\odot}$ which noise effect is not too much), the BT models boost the normalized radial number density near the center by a factor of two.
As suggested by Springel et al. (2008), within the range of $10^{5}~{}{\rm M}_{\odot}\sim 10^{10}~{}{\rm M}_{\odot}$, massive subhalos tend to exist in the radially outer region while different mass bins follow the same line. Both our BT and PL simulation results also show this trend. Although the BT subhalo number density is higher, the radial range where subhalos exist within that mass range does not change.
Figure 8: The subhalo radial number density profiles at four different mass bins for m12i (colored solid) and m12f (colored dashed). In each subfigure (also in each mass bin): The upper panel shows the radial number density $n(r)$ normalized by $n\right>$ in Equation 11, for the PL (orange), the BT_deep (green), and the BT_soft (red) models. The grey solid line shows the equivalent substructure profile from Figure 12 of Ref. Springel et al. (2008). The bottom panel shows the ratios of the BT normalized number density to the PL normalized number density.
IV.2.5 Cumulative Substructure Mass Fraction
In Figure 9, we present the cumulative substructure mass fraction: the mass of particles belonging to subhalos over the total mass, within a certain radius to the main halo center $r$. This substructure fraction could be useful for future strong lensing studies Vegetti and Koopmans (2009) on the BT models. We take the cumulative substructure mass fraction function for the PL LCDM model from Ref. Lovell et al. (2014) as the Lovell CDM result. Additionally, we rescale the radius by $R_{200c}$ of the main halo, to eliminate the influence of the different main halo sizes.
The upper panel of Figure 9 shows a good match between the Lovell CDM result and our PL simulation for m12i. But, unexpectedly, the PL simulation of m12f is higher than the Lovell CDM result. This is because, in m12f, there is a subhalo with $7.7\times 10^{8}~{}{\rm M}_{\odot}$ within $0.1R_{200c}$, which can account for the unusually high substructure fraction for the m12f PL simulation.
We find that the BT_deep and BT_soft lines are higher than the Lovell CDM result and our BT simulations, implying that there are more substructures with the BT than PL PPS. The enhancement is especially strong in the inner region, where strong lensing is likely to probe.
To quantify the difference, in the bottom panel, we show the ratio of the cumulative substructure mass fraction between the BT and PL simulations. For BT_deep, the fraction is enhanced by a factor of ten at the inner region ($0.1R_{200c}$). As the radius increases, the enhancement ratio drops and finally becomes stable around a factor of two at the outer region. This is also consistent with our result in subsubsection IV.2.2: the BT models can double the total mass residing in resolved subhalos.
Figure 9: The cumulative mass fraction in substructures as a function of the radius for PL (orange), BT_deep (green), and BT_soft (red). The upper panel shows the cumulative substructure fraction function, with $r$ (the subhalo distance from the main-halo center) scaled by $R_{200c}$ of the main halo. The grey solid line shows the equivalent substructure profile from Ref. Lovell et al. (2014). The bottom panel shows the ratios of the BT substructure fractions to the PL substructure fractions.
IV.2.6 Subhalo $R_{\rm max}-V_{\rm max}$ relationship
We present the subhalo $R_{\rm max}-V_{\rm max}$ relationship in Figure 10 for all the subhalos within $R_{200c}$ from the main halo center, with m12i and m12f, respectively. It is a measurement of the subhalo central density, thus astrophysically interesting: for a specific $V_{\rm max}$ bin, a larger(smaller) $R_{\rm max}$ usually means a lower(higher) central density. The black line is taken from Ref. Grand and White (2021) (Robert) for the PL LCDM cosmology, considering all the subhalos within $R_{200c}$ of the main halos.
The upper panel of Figure 10 shows a good match between the Robert result and our PL simulation, for both m12i and m12f. The two BT models have much smaller $R_{\rm max}$ than the PL model. The BT_soft model has a lower $R_{\rm max}$ than the BT_deep model. This implies while subhalos in BT_deep are more concentrated than in PL, the BT_soft model is even more concentrated than the BT_deep model. The unresolved regions for m12i PL and m12f PL are indicated with shaded areas, following the same definition as in subsubsection IV.2.3.
In the bottom panels, the BT models suppress $R_{\rm max}$ at nearly the whole resolved $V_{\rm max}$ range. The effects in m12i and m12f are similar: for the BT_deep model, $R_{\rm max}$ decreases by $\sim 40\%$; for the BT_soft model, $R_{\rm max}$ decreases nearly by half.
Figure 10: The subhalo $V_{\rm max}$-$R_{\rm max}$ relations for m12i (left figure) and m12f (right figure) respectively. For both figures:
The upper panel shows the subhalo $R_{\rm max}-V_{\rm max}$ relations for all the subhalos within $R_{200c}$ from the main halo center with the PL (orange), BT_deep (green), and BT_soft (red) models. The colored solid lines depict the median relation found by binning the subhalos according to their $V_{\rm max}$ values. The colored regions surrounding the colored lines illustrate the 16th-84th percentiles around the medians. The black solid line is the result of Ref. Grand and White (2021) for the PL LCDM model, taking account of all the subhalos within $R_{200c}$ of the main halo. The bottom panel shows the ratios of $R_{\rm max}$ between the BT and PL models. The shaded area is the same as that in Figure 7.
V CONCLUSIONS
The standard model of cosmology – the Power Law Primordial Power Spectrum + the Lambda Cold Dark Matter – has been successful over the past several decades, especially on the large scales of the universe Planck Collaboration et al. (2020a, b); Troxel et al. (2018); Blanton et al. (2017); Chabanier et al. (2019). However, there are still possible tensions between the observational results and the theoretical simulations.
JWST’s high redshift observations revealed that galaxies could form earlier than those in the standard model of cosmology Labbé et al. (2023); Boylan-Kolchin (2023); Lovell et al. (2023). The strong lensing observations prefer more substructures Macciò and Miranda (2006); Xu et al. (2009, 2015). Furthermore, the recent observations Müller et al. (2024); Homma et al. (2024) and theories Kim et al. (2018) suggested there could be a “too many satellites” problem in the nearby universe. These motivate us to consider alternative cosmological models with an enhanced Primordial Power Spectrum at small scales ($>1~{}{\rm Mpc}^{-1}$), namely the blue-tilted (BT) model.
In this study, we have conducted cosmological zoom-in simulations of MW host-size halos with the blue-tilted primordial power spectrum. We have considered two blue-tilted models from Ref. Hirano and Yoshida (2024): BT_soft with the enhancement at $\gtrsim 1~{}{\rm Mpc}^{-1}$ and BT_deep with the enhancement at $\gtrsim 4~{}{\rm Mpc}^{-1}$. They are all within the accepted parameter space according to the JWST observation Hirano and Yoshida (2024) and the dwarf galaxy central density observation Dekker and Kravtsov (2024).
Our main results are summarized in the following:
  1. 1.
    We have found that BT_soft can enhance the mass, radius, $V_{\rm max}$, concentration (Table 2), and the radial density profile (subsubsection IV.2.1) of the main halo. However, the effects of the BT_deep model on the main halo are much smaller, since it is blue-tilted at a smaller scale in the primordial power spectrum, which corresponds to a smaller mass scale (Equation 5).
  2. 2.
    The BT models enhance the subhalo mass function by $\sim 2$ and the subhalo $V_{\rm max}$ function by $\gtrsim 3$ for a wide range of halo masses ($\sim 10^{6}~{}{\rm M}_{\odot}-10^{9}~{}{\rm M}_{\odot}$) and maximum circular velocities ($\sim 5~{}{\rm km/s}-50~{}{\rm km/s}$) (subsubsection IV.2.2 and subsubsection IV.2.3). The enhancement ratios, defined as the numbers of subhalos in BT over that in PL, follow inverse S shape functions (Equation 8) in both mass and $V_{\rm max}$ functions.
    In both the BT_deep and BT_soft simulations, the enhancements in the subhalo mass function are similar. However, the BT_soft simulation has a stronger enhancement in the subhalo $V_{\rm max}$ function than the BT_deep simulation.
  3. 3.
    The BT primordial power spectrum increases the number density of subhalos in the inner region of the main host halo by more than a factor of two, compared to PL (subsubsection IV.2.4). In the BT simulations, the number densities of subhalos in the lower mass bins are more centrally concentrated.
  4. 4.
    We have found that the BT models can boost the mass fraction of the substructures in the inner region ($<0.1R_{200c}$) by an order of magnitude (subsubsection IV.2.5). Even around $R_{200c}$, the BT models could have a two-fold increase in the substructure mass fraction, compared to PL.
  5. 5.
    At a fixed $V_{\rm max}$, the BT models reduce the mean $R_{\rm max}$ values of the subhalos by $40\%-50\%$ (subsubsection IV.2.6). This reveals that subhalos are more concentrated in BT than PL.
To constrain the primordial power spectrum with our simulations, we require comparisons with observations, including the statistics of nearby satellite galaxies and the strong lensing observations of Milky-Way size host halos. However, at these small scales, baryonic physics cannot be ignored, e.g. the tidal stripping of subhalos with baryonic disks (Garrison-Kimmel et al., 2017), the uncertainties in the stellar-to-halo mass relation (Sales et al., 2022), the baryonic effects on the inner density profiles (Chan et al., 2015), and the lensing effects of the stellar structures (Gilman et al., 2017; Hsueh et al., 2016). Therefore, it is necessary to incorporate regulator-type models (Finlator and Davé, 2008; Kravtsov and Manwadkar, 2022), semi-analytic models (Cole et al., 2000; Benson, 2012), or even cosmological galaxy simulations (Hopkins et al., 2018; Schaye et al., 2023) to disentangle the effects of the primordial power spectrum from baryons.
Acknowledgments
We thank Matthieu Schaller, Yiming Zhong, Ariane Dekker, and Kin-Wang Ng for their thoughtful comments. The research in this paper used the SWIFT open-source simulation code (http://www.swiftsim.com, Schaller et al. (2018)) version 1.0.0. We also used SWIFTSIMIO (Borrow and Borrisov, 2020; Borrow and Kelly, 2021) to generate the projection maps. This work also made use of matplotlib (Hunter, 2007), numpy (van der Walt et al., 2011), scipy (Jones et al., 2001; Virtanen et al., 2020) and NASA’s Astrophysics Data System. We acknowledge the support of the CUHK Central High-Performance Computing Cluster, on which the computation in this work has been performed; and we thank Edward So, Anny Cheung, and Nicky Leung for their assistance.
TKC is supported by the ‘Improvement on Competitiveness in Hiring New Faculties’ Funding Scheme from the Chinese University of Hong Kong (4937210, 4937211, 4937212). TKC and JW are partially supported by the Direct Grant project (4053662) from the Chinese University of Hong Kong. VJFM acknowledges support by NWO through the Dark Universe Science Collaboration (OCENW.XL21.XL21.025).
Appendix A Results from another halo finder: VELOCIraptor
Apart from HBT+HERONS, we also identify the halos and subhalos with VELOCIraptor (VR) (Elahi et al., 2019), a code based on the friends-of-friends (FOF) algorithm that identifies peaks in 6D phase-space to find subhalos (both the configuration space and phase space). We compare the HBT+HERONS and VR results on the subhalo mass function, subhalo $V_{\rm max}$ function, and subhalo radial distribution in subsection A.1.
Currently, SOAP (see subsection III.3) only allows the calculation of spherical overdensity properties for field halos. Besides, HBT+HERONS (without SOAP post process) could only calculate limited kinds of SO properties. Hence, we used VR (without SOAP post process) to calculate the virial masses for all the halos (also including subhalos) and show the $M_{\rm vir}-V_{\rm max}$ relationship in subsection A.2. Usually, the spherical overdensity properties will include all of the particles within the SO radius, regardless of whether they are bound to the main halo. However, for subhalos, VR only considers the bound particles [^8] See Mass and radius properties in https://velociraptor-stf.readthedocs.io/en/latest/output.html. Otherwise, the subhalos near the center of the main halo will have unphysically large masses.
Our essential parameters for VR are listed in footnote 11, based on the sample configuration for 6D dark matter-only simulation  [^9] https://velociraptor-stf.readthedocs.io/en/latest/_downloads/97ba16973bcff9d4afe940c2e13d12d2/sample_dmcosmological_6dfof_subhalo.cfg. We set the most bound particle’s position as the reference frame center to be consistent with HBT+HERONS for comparison in subsection A.1.
Parameters Value
FoF_Field_search_type 3
Minimum_halo_size 20
Bound_halos 1
Kinetic_reference_frame_type 0
Reference_frame_for_properties 1
Table 3: Selected parameter settings in the VELOCIraptor configuration file. For detailed explanations of these parameters, refer to the usage section in the documentation  [^11] https://velociraptor-stf.readthedocs.io/en/latest/usage.html.
A.1 Comparison between halo finders: HBT+HERONS and VELOCIraptor
Here, we choose m12i to analyze the difference between the two halo finders. VR stands for VELOCIraptor in both the text and figures. In this subsection, we also use SOAP to process VR and HBT+HERONS catalogs in a consistent manner.
However, for the previous literatures we fit with, COCO (Hellwing et al., 2016) and Aquarius (Springel et al., 2008) use SUBFIND (Springel et al., 2001) as their halo finder, which is a configuration space finder; Cautun fit line (Cautun et al. (2014)) is derived via ROCKSTAR (Behroozi et al., 2013), a phase space finder. So there would be inherent offset simply due to the choice of halo finder, for the fitting lines. As a result, the deviation from fitting lines could not imply the capability of the halo finder. The relative ratio between the two halo finders would be more meaningful.
Figure 11: The cumulative scaled subhalo mass functions for m12i, with two halo finders HBT+HERONS (colored solid) and VR (colored dashed), respectively. The upper panel shows the subhalo mass function for the PL (orange), BT_deep (green), and BT_soft (red) models, with the subhalo mass $M_{\rm sub}$ scaled by $M_{200c}$ of the main halo. The bottom panel shows the ratios of the VR to HBT+HERONS number. The shaded areas are the same as that in Figure 6. The shaded areas overlap because we set the minimum size of the subhalo to be the same in both halo finders.
The subhalo mass function comparison between VR and HBT+HERONS is shown in Figure 11. These two halo finders agree with each other for $10^{-5}<\mu<10^{-3},10^{7}~{}{\rm M}_{\odot}<M_{\rm sub}<10^{9}~{}{\rm M}_{\odot}$. However, HBT+HERONS is higher and closer to the COCO fitting line than VR at the high mass end, implying that HBT+HERONS is better at finding more massive subhalos. But at the low mass end, VR appears to have more subhalos than HBT+HERONS.
Figure 12: The cumulative scaled subhalo $V_{\rm max}$ functions for m12i, with two halo finders HBT+HERONS (colored solid) and VR (colored dashed), respectively. The upper panel shows the subhalo $V_{\rm max}$ function for the PL (orange), BT_deep (green), and BT_soft (red) models, with the subhalo $V_{\rm max}$ scaled by $V_{200c}$ of the main halo. The bottom panel shows the ratios of the VR to HBT+HERONS number. The shaded areas are the same as those in Figure 7.
The subhalo $V_{\rm max}$ function comparison between VR and HBT+HERONS is shown in Figure 12. These two halo finders give similar results except at the low $V_{\rm max}$ end near the resolution limit. When approaching the limit $V_{\rm max}$ value, VR deviates from the power-law behavior later than HBT+HERONS.
Figure 13: The subhalo radial number density profiles at four different mass bins for m12i, using different halo finders HBT+HERONS (colored solid) and VR (colored dashed). We consider the PL (orange), BT_deep (green), and BT_soft (red) models in each subfigure (also in each mass bin): the upper panel is the same definition as that in Figure 8; the bottom panel shows the ratios of the VR results to the HBT+HERONS results.
The subhalo radial distribution comparison between VR and HBT+HERONS is shown in Figure 13. In mass bins larger than $10^{7}~{}{\rm M}_{\odot}$, VR and HBT+HERONS have nearly the same results. But, at the lower mass end, VR has much fewer subhalos in the inner region than HBT+HERONS and the Aquarius result.
To summarize, VR and HBT+HERONS behave similarly for the subhalo mass and $V_{\rm max}$ functions. However, for the radial properties (the subhalo radial distribution function), VR gaves a lower halo count near the host halo center for the low halo mass bin.
A.2 Halo $M_{\rm vir}-V_{\rm max}$ relationship
Figure 14: The subhalo $M_{\rm vir}-V_{\rm max}$ relations for m12i (left figure) and m12f (right figure) respectively. For both figures:
The upper panel shows the $M_{\rm vir}-V_{\rm max}$ relationship for all the field halos and subhalos within the simulation domain, with the PL (orange), BT_deep (green), and BT_soft (red) models. The colored solid lines depict the median relation by binning the subhalos according to their $V_{\rm max}$ values. The colored regions surrounding the colored lines illustrate the 16th-84th percentiles around the medians. The black solid line results from Ref. Rodríguez-Puebla et al. (2016) for the PL LCDM model, including both the field halos and subhalos. The bottom panel shows the ratios of BT $M_{\rm vir}$ to PL $M_{\rm vir}$. The shaded area is the same as that in Figure 7.
We present the halo $M_{\rm vir}-V_{\rm max}$ relationship in Figure 14 for all the field halos and subhalos. The black line is the extrapolation of Ref. (Rodríguez-Puebla et al., 2016) for the PL LCDM cosmology, including both the field halos and subhalos.
The upper panels of Figure 14 show a good match between the BolshoiP & MDPL simulation and our PL simulation, for both m12i and m12f. The BT lines are lower than the PL lines, implying the BT halos are more concentrated. We also observe the BT_soft model has a lower $M_{\rm vir}$ than the BT_deep model for the same $V_{\rm max}$.
In the bottom panels, we show that the BT models suppress $M_{\rm vir}$ at the resolved $V_{\rm max}$ range. The suppressions in m12i and m12f are similar: for the BT_deep model, $M_{\rm vir}$ decreases by around 30%; for the BT_soft model, $M_{\rm vir}$ decreases by around 50%. At high velocities, the BT models are closer to the PL model’s $M_{\rm vir}$ value. This implies the BT effects are stronger in the lower mass halos.
Appendix B Resolution limit study
To investigate whether the conclusion still holds at a higher resolution, we take m12i as an example to study the consistency between different resolutions. In the high-resolution run, we reduce the particle mass to one-eighth of the original (and boost the number of particles by a factor of 8 times). The main halo properties for the high-resolution runs are also shown in Table 2.
Figure 15: The cumulative scaled subhalo mass functions for m12i, under fiducial resolution (colored solid) and high resolution (colored dashed), respectively.
The upper panel shows the subhalo mass function for the PL (orange), BT_deep (green), and BT_soft (red) models, with the subhalo mass $M_{\rm sub}$ scaled by $M_{200c}$ of the main halo. The bottom panel shows the ratios of high-resolution to fiducial-resolution numbers. The shaded areas are the same definition as that in Figure 6. However, the high-resolution run has a smaller particle mass, so the lower bound is also lower.
The subhalo mass function comparison between the high and fiducial resolution is shown in Figure 15. From the upper panel, we find that the high-resolution run can extend the power-law behavior to a lower mass range, matching the COCO fitting line. We find the overlapping of the deep and soft models in Figure 6 still holds at a higher resolution. From the bottom panel, by comparing the high and fiducial runs, we find that the fiducial run deviates from the high-resolution run at around $\mu\sim 10^{-5.5}$. This implies that we can trust the results from halos with more than 50 particles. The main conclusions in Figure 6 still hold.
Figure 16: The cumulative scaled subhalo $V_{\rm max}$ functions for m12i, under fiducial resolution (colored solid) and high resolution (colored dashed), respectively.
The upper panel shows the subhalo $V_{\rm max}$ function for the PL (orange), BT_deep (green), and BT_soft (red) models, with $V_{\rm max}$ of subhalo scaled by $V_{200c}$ of the main halo. The bottom panel shows the ratios of the high-resolution to fiducial-resolution numbers. The shaded areas are the same definition as that in Figure 7. Similar to Figure 15, the high-resolution run can resolve the subhalos with a lower $V_{\rm max}$.
The subhalo $V_{\rm max}$ function comparison between the high and fiducial resolution is shown in Figure 16. Similar to Figure 15, we find the power-law behavior to a lower $V_{\rm max}$ range. The BT_deep line is higher than the BT_soft line, which is higher than the PL line. Thus, as in subsubsection IV.2.3, we find that the subhalo $V_{\rm max}$ function is a better distinguisher for the BT models than the subhalo mass function. From the bottom panel, we find that the $V_{\rm max}$ of subhalo with 100 particles is reliable (with 30%).
Figure 17: This figure is the same as the upper panel of Figure 10, but with a higher resolution. The upper panel shows the subhalo $R_{\rm max}-V_{\rm max}$ relationship with the PL (orange), BT_deep (green), and BT_soft (red) models, in the high-resolution run of m12i; the colored solid lines depict the median relation found by binning the subhalos according to their $V_{\rm max}$ values. The colored regions surrounding the colored lines illustrate the 16th-84th percentiles around the medians. The black solid line is the result of Ref. Grand and White (2021) for the PL LCDM model, taking account of all the subhalos within $R_{200c}$ of the main halo. The bottom panel shows the ratios of $R_{\rm max}$ between the BT models and the PL model. The shaded area is the same definition as in Figure 16.
The subhalo $R_{\rm max}-V_{\rm max}$ relationship at a higher resolution for m12i is shown in Figure 17. From the upper panel, we find that the PL result starts to diverge with the reference line Grand and White (2021) from $V_{\rm max}\lesssim 10~{}{\rm km/s}$. It suggests that the $R_{\rm max}$ of subhalos at the low- $V_{\rm max}$ end is overestimated in the fiducial-resolution run (see the upper panels of Figure 10). However, for the subhalos whose $V_{\rm max}$ is beyond $10~{}{\rm km/s}$, the results on different resolutions are consistent; besides, the relative differences among different PPS models still hold: from the bottom panel, there is a $40\%$ ($50\%$) decrease in BT_deep (BT_soft) in $R_{\rm max}$ at the same $V_{\rm max}$ bin.
Figure 18: The subhalo radial number density profiles at four different mass bins for m12i, with the fiducial (colored solid) and high resolution (colored dashed), respectively. We consider the PL (orange), BT_deep (green), and BT_soft (red) models in each subfigure (also in each mass bin): The upper panel is the same definition as that in Figure 8. The bottom panel shows the ratios of the result under the high resolution to the result with the fiducial resolution.
The subhalo radial distribution comparison between high resolution and fiducial resolution is shown in Figure 18. The upper panels show that the relative differences among different PPS models still hold: at the inner region $BT\_soft>BT\_deep>PL$. The bottom panels show that for the mass bins larger than $10^{7}~{}{\rm M}_{\odot}$, the result is quite close at these two resolutions within scatters. However, consistent with what we see in the mass function comparison, the mass bin $10^{6}~{}{\rm M}_{\odot}<M_{\rm sub}<10^{7}~{}{\rm M}_{\odot}$, part of which is close to the minimum halo size of the fiducial-resolution run, shows the largest difference at two resolutions: the fiducial case has over-predicted subhalo number by a quarter in the inner region, while underestimated that by around 30% in the outer region.
Summary: We have proven the fiducial-resolution simulation’s conclusions still hold at a higher resolution. We can trust results from halos with $\gtrsim 50$ particles for the subhalo mass function and $V_{\rm max}\gtrsim 10km/s$ for the subhalo $R_{\rm max}-V_{\rm max}$ relation and the subhalo $V_{\rm max}$ function.

### v2

@@
-Cosmological Zoom-In Simulations of Milky Way Host Size Dark Matter Halos with a Blue-Tilted Primordial Power Spectrum
+Cosmological Zoom-In Simulations of Milky Way Host Mass Dark Matter Halos with a Blue-Tilted Primordial Power Spectrum
 Abstract
-Recent observations from the James Webb Space Telescope revealed a surprisingly large number of galaxies formed at high redshift.
-Along with strong lensing studies and nearby galaxy observations, these could challenge the standard Lambda Cold Dark Matter cosmology with a power-law primordial power spectrum.
-In this study, we conduct high-resolution cosmological zoom-in dark matter-only simulations of Milky Way host size halos with a blue, tilted primordial power spectrum ($P(k)\propto k^{m_{s}}$ with $m_{s}>1$ at small scales $>1~{}{\rm Mpc}^{-1}$).
-We find that the blue-tilted subhalo mass functions can be enhanced by more than a factor of two for subhalo masses $M_{\rm sub}\lesssim 10^{10}~{}{\rm M}_{\odot}$, whereas the subhalo $V_{\rm max}$ functions can be enhanced by a factor of four for maximum circular velocities $V_{\rm max}\lesssim 30~{}{\rm km/s}$.
+Recent observations from the James Webb Space Telescope revealed a surprisingly large number of galaxies at high redshift, challenging the standard Lambda Cold Dark Matter cosmology with a power-law primordial power spectrum.
+Previous studies alleviated this tension with a blue tilted primordial power spectrum ($P(k)\propto k^{m_{s}}$ with $m_{s}>1$ at small scales $>1~{}{\rm cMpc}^{-1}$).
+In this study, we examine whether the blue tilted model can boost dark matter substructures especially at low redshift, thereby addressing other potential challenges to the standard cosmology.
+First, substructures in the standard cosmological model may not be sufficient to explain the anomalous flux ratio problem observed in strong gravitational lensing.
+Second, the number of observed nearby satellite galaxies could be higher than the theoretical predictions of the standard cosmology, after completeness correction and tidal stripping by baryonic disks.
+To study the impact of a blue tilted primordial power spectrum on substructures, we perform high-resolution cosmological zoom-in dark matter-only simulations of Milky Way host mass halos, evolving to redshift $z=0$.
+At $z=0$, we find that the blue-tilted subhalo mass functions can be enhanced by more than a factor of two for subhalo masses $M_{\rm sub}\lesssim 10^{10}~{}{\rm M}_{\odot}$, whereas the subhalo $V_{\rm max}$ functions can be enhanced by a factor of four for maximum circular velocities $V_{\rm max}\lesssim 30~{}{\rm km/s}$.
 The blue-tilted scaled cumulative substructure fraction can be an order of magnitude higher at $\sim$10of the virial radius.
@@
 (2019).
-However, there are relatively sparse constraints for the scales smaller than $\sim 1~{}{\rm Mpc}^{-1}$.
+However, there are relatively sparse constraints for the scales smaller than $\sim 1~{}{\rm cMpc}^{-1}$.
 In particular, there have been active debates over the so-called small-scale crisis of LCDM Bullock and Boylan-Kolchin (2017), including the Missing Satellite Problem (MSP, (Klypin et al., 1999; Moore et al., 1999)), the Core-Cuspy problem (Salucci and Burkert, 2000; Flores and Primack, 1994), and the Too-Big-To-Fail problem (Boylan-Kolchin et al., 2011), etc.
 This leaves room for other models beyond the standard cosmology, including different models of dark matter, effects of baryons, and modifications to the PPS.
-One possible modification is to enhance the PPS on small scales (${\rm>1~{}{\rm Mpc}^{-1}}$) (Covi and Lyth, 1999; Martin and Brandenberger, 2001; Gong and Sasaki, 2011), deviating from the PL PPS.
+One possible modification is to enhance the PPS on small scales (${\rm>1~{}{\rm cMpc}^{-1}}$) (Covi and Lyth, 1999; Martin and Brandenberger, 2001; Gong and Sasaki, 2011), deviating from the PL PPS.
 In these small-scale enhanced models, PPS will follow the traditional PL model on large scales, preserving the success of the PL PPS, while having an enhancement on small scales.
@@
 Hirano et al.
-(2015) considered blue-tilted (BT) models, which are broken power law PPSs with a higher power law index in scales $>1~{}{\rm Mpc}^{-1}$.
-They found a significant boost in dark matter halos with BT BBS, especially at high redshift.
-More complicated forms also exist, such as several bumps in the PPS Tkachev et al.
+(2015) considered blue-tilted (BT) models, which are broken power law PPSs with a higher power law index in scales $>1~{}{\rm cMpc}^{-1}$.
+They found a significant boost in the number of dark matter halos with BT PPS, especially at high redshift.
+Hence, using the BT model, Ref.
+Hirano et al.
+(2015) explained the large number of high redshift massive galaxies observed by the James Webb Space Telescope (JWST) Gardner et al.
+(2006), which is unexpected from the PL model.
+More complicated forms of enhancements also exist, such as several bumps in the PPS Tkachev et al.
 (2024).
-With the recently launched James Webb Space Telescope (JWST) Gardner et al.
-(2006), Ref.
-Labbé et al.
-(2023) found a large number of galaxies formed in the high redshift universe ($z>10$), which may be in tension with the standard cosmology Boylan-Kolchin (2023); Lovell et al.
-(2023).
-Ref.
-Parashari and Laha (2023) eased this tension with the BT formalism, based on the Press-Schechter formalism Press and Schechter (1974) and the Sheth-Tormen correction Sheth and Tormen (1999).
-Ref.
-Hirano and Yoshida (2024) performed cosmological N-body simulations with the BT model, achieving similar conclusions that the BT model could match the JWST observations without invoking ultra-high star formation rates.
-However, other mechanisms in galaxy formation models (e.g.
-Refs.
-Dekel et al.
-(2023); Haslbauer et al.
-(2022); Yung et al.
-(2024)) could potentially explain the JWST data.
-Thus, apart from the high redshift massive galaxies, additional observations are required to distinguish the BT and PL models.
-The strong lensing observations can provide another constraint.
-The asymptotic flux ratio relation, which serves as a general principle in smooth lensing potentials (Mao, 1992), is often observed to be violated.
-This is taken as evidence for the presence of the substructures around lensing galaxies Mao and Schneider (1998).
-However, in the PL LCDM cosmology, there could be not enough low mass substructures in lensing galaxies to explain the violations Macciò and Miranda (2006); Xu et al.
-(2009, 2015).
-This could suggest extra structures not present in dark matter-only simulations under the standard PL LCDM cosmology.
+Instead of focusing on high redshift, in this study, we examine the impact of the BT model on the low redshift dark matter substructures.
+There are two major motivations for this: (1) the anomalous flux ratio problem in strong gravitational lensing and (2) potentially the too many satellite problem.
+The strong gravitational lensing observations can probe invisible substructures, so it can constrain dark matter substructures and, thus, the cosmological model.
+For example, the asymptotic flux ratio relation, which serves as a general principle in smooth lensing potentials (Mao, 1992), is often observed to be violated.
+This is taken as evidence for the presence of the substructures around lensing galaxies (Mao and Schneider, 1998).
+However, in the PL cosmology, there could be not enough low mass substructures in lensing galaxies to explain the violations (Macciò and Miranda, 2006; Xu et al., 2009, 2015).
+This could suggest extra structures not present in dark matter-only simulations under the standard PL cosmology.
 On the other hand, Ref.
@@
 (2022) constrained sub-galactic PPS to be consistent with PL with the 11 strongly lensed and quadruply image quasars, albeit with possibilities with blue/red tilted PPS on small scales.
-The density profiles and statistics of nearby low-mass dwarf galaxies can also inform us about the nature of dark matter, baryonic physics, and the PPS.
-For example, the observed central densities of dwarf galaxies can provide insights into the nature of dark matter Rocha et al.
-(2013); Yang et al.
-(2024) or constraining baryonic physics Di Cintio et al.
-(2014); Chan et al.
-(2015); Tollet et al.
-(2016); Lazar et al.
-(2020).
-The “missing satellite problem” describes the discrepancy between the number of subhalos from cosmology dark matter simulations and the observed satellite galaxy population (Klypin et al., 1999; Moore et al., 1999).
-This problem could signify a challenge to the standard PL LCDM cosmology, or the importance of including accurate galaxy formation physics in dwarf galaxies.
-Nearby dwarf galaxies also provide constraints on the PPS.
-For example, Refs.
-Esteban et al.
-(2024); Dekker and Kravtsov (2024) employed semi-analytic models to constrain the BT PPS parameter space, using the observed relation between half-light radius and V-band luminosity function in nearby dwarf galaxies.
+Another motivation for the BT model is that the number of satellite galaxies could be higher than the expected from the PL model.
 Thanks to the recent improvements in observations, many more ($\gtrsim 50$) faint satellite galaxies have been discovered in Milky Way (MW) and nearby MW mass hosts Drlica-Wagner et al.
@@
 Kim et al.
-(2018) argued that the “missing satellites problem” can be solved.
-It could even become the “too many satellites problem”, meaning that there are too many satellite galaxies inferred from observations than expected in simulations/models Kim et al.
-(2018).
-This happens if some mechanisms can reduce the number of satellites in simulations /models.
+(2018) argued that there could be too many satellite galaxies inferred from observations than expected in simulations/models.
+This happens if some mechanisms can reduce the number of satellite galaxies in simulations / models.
 For example, a baryonic disk can tidally strip subhalos Garrison-Kimmel et al.
 (2017), and/or the minimum subhalo mass to host an observable galaxy is large Graus et al.
+(2019); Jethwa et al.
 (2018).
 This “too many satellites problem” motivates us to explore mechanisms to boost the number of subhalos, which includes the BT model.
-Motivated by high redshift, strong lensing, and nearby galaxy observations, we explore the impact of the BT PPS on the substructures in dark matter halos, which, to our knowledge, has not been previously investigated with numerical simulations.
-We perform cosmological zoom-in dark matter-only simulations of MW host size halos with BT PPS to study its substructures at $z=0$.
+A related constraint on the primordial power spectrum, though not the focus of this paper, is the density profiles of nearby low-mass dwarf galaxies.
+For example, the observed central densities of dwarf galaxies can provide insights into the nature of dark matter Rocha et al.
+(2013); Yang et al.
+(2024) or constraining baryonic physics Di Cintio et al.
+(2014); Chan et al.
+(2015); Tollet et al.
+(2016); Lazar et al.
+(2020).
+Refs.
+Esteban et al.
+(2024); Dekker and Kravtsov (2024) employed semi-analytic models to constrain the BT PPS parameter space, using the observed relation between half-light radius and V-band luminosity function in nearby dwarf galaxies.
+Motivated by strong lensing, and nearby galaxy observations, we explore the impact of the BT PPS on the substructures in dark matter halos at low redshift, which, to our knowledge, has not been previously investigated with numerical simulations.
+We perform cosmological zoom-in dark matter-only simulations of MW host mass halos with BT PPS to study its substructures at $z=0$.
 Note that we only modify the PPS but still consider only cold dark matter (i.e.
@@
 We present our main results in section IV and conclude in section V.
-Additionally, as the first work to use HBT-HERONS (see subsection III.3) for zoom-in simulations, we show the comparison between HBT+HERONS and another halo finder VELOCIraptor in Appendix A.
+Additionally, as the first work to use HBT-HERONS (see subsection III.3) for zoom-in simulations, we show the comparison between HBT-HERONS and another halo finder VELOCIraptor in Appendix A.
 We also run higher-resolution simulations and compare the fiducial results in Appendix B.
 II Theoretical Overview
-Figure 1: The conceptual steps of the cosmological stages and our simulation steps.
-The upper panel shows the universe’s chronology, along with redshifts of some critical eras; the lower panel shows the numerical tools we used.
 The upper panel of Figure 1 shows the cosmic chronology.
@@
 In this paper, along with the PL model, we pick two BT models from Ref.
-Hirano and Yoshida (2024) [^1] For $k_{p}$, the value in our table is a float number to several digits because original models use $h~{}\rm{Mpc}^{-1}$ as a unit of $k_{p}$, but we convert it into ${\rm Mpc}^{-1}$.: BT_deep and BT_soft, listed in Table 1.
+Hirano and Yoshida (2024) [^1] For $k_{p}$, the value in our table is a float number to several digits because original models use $h~{}\rm{cMpc}^{-1}$ as a unit of $k_{p}$, but we convert it into ${\rm cMpc}^{-1}$.: BT_deep and BT_soft, listed in Table 1.
 BT_soft with a reasonable star formation efficiency $\sim 0.2$ can explain the large number of galaxies formed at high redshift as observed by JWST Hirano and Yoshida (2024).
@@
 (2022).
+Figure 1: The conceptual steps of the cosmological stages and our simulation steps.
+The upper panel shows the universe’s chronology, along with redshifts of some critical eras; the lower panel shows the numerical tools we used.
 For all of our PPS models, the matter power spectra at z=1089 and the parameter settings are shown in Figure 2 and Table 1 respectively.
-BT_soft deviates from PL at $0.7~{}{\rm Mpc}^{-1}$, whereas BT_deep deviates at $3.51~{}{\rm Mpc}^{-1}$, but they have the same power law index afterward.
+BT_soft deviates from PL at $0.7~{}{\rm cMpc}^{-1}$, whereas BT_deep deviates at $3.51~{}{\rm cMpc}^{-1}$, but they have the same power law index afterward.
+Models Related parameters
+PL Power Law Primordial Power Spectrum
+$n_{s}=0.961$
+BT_deep $k_{p}=3.51\ \rm{cMpc}^{-1}$ $m_{s}=1.5$
+BT_soft $k_{p}=0.702\ \rm{cMpc}^{-1}$ $m_{s}=1.5$
+Table 1: The parameters of all the chosen models.
+$k_{p}$ is the wave vector at which the BT PPS would deviate from the PL PPS.
+$m_{s}$ is the enhanced spectral index for $k>k_{p}$, at the small scales.
+For other cosmological parameters, see subsubsection III.2.1.
 Figure 2: The upper panel shows the power spectra for the matter density perturbation at z=1089 with PL (orange), BT_deep (green), and BT_soft (red).
@@
 The model parameters are listed in Table 1.
-Models Related parameters
-PL Power Law Primordial Power Spectrum
-$n_{s}=0.961$
-BT_deep $k_{p}=3.51\ \rm{Mpc}^{-1}$ $m_{s}=1.5$
-BT_soft $k_{p}=0.702\ \rm{Mpc}^{-1}$ $m_{s}=1.5$
-Table 1: The parameters of all the chosen models.
-$k_{p}$ is the wave vector at which the BT PPS would deviate from the PL PPS.
-$m_{s}$ is the enhanced spectral index for $k>k_{p}$, at the small scales.
-For other cosmological parameters, see subsubsection III.2.1.
-The wave number $k$ can be connected with a mass scale $M_{l}$ at z=0 Bullock and Boylan-Kolchin (2017), by assuming the sphere radius $r_{l}$ to be half wavelength $\frac{\pi}{k}$:
-$M_{l}=$ $\frac{4\pi}{3}{r_{l}}^{3}\rho_{m}=\frac{\Omega_{m}H_{0}^{2}}{2G}r_{l}^{3}$
-$=$ $1.71\times 10^{11}\left(\frac{\Omega_{m}}{0.3}\right)\left(\frac{H_{0}}{70}\right)^{2}\left(\frac{r_{l}}{1~{}\rm{Mpc}}\right)^{3}~{}{\rm M}_{\odot}.$ (4)
-We further give the pivot mass $M_{p}$ under which the BT model can enhance the structure formation:
-$M_{p}=5.29\times 10^{12}\left(\frac{\Omega_{m}}{0.3}\right)\left(\frac{H_{0}}{70}\right)^{2}\left(\frac{k_{p}}{1~{}{\rm Mpc}^{-1}}\right)^{-3}~{}{\rm M}_{\odot}.$ (5)
+Figure 3: The cumulative halo mass function generated by genmf, at z=9 (dashed line) and z=0 (solid line), with PL (orange), BT_deep (green), and BT_soft (red).
+The upper panel shows the cumulative halo mass function per unit comoving volume.
+The bottom panel shows the ratios of BT to PL.
+The triangle symbols on the solid lines show the pivot masses $M_{p}$ for BT_deep (green) and BT_soft (red), respectively.
+We can associate a comoving wavenumber $k$ with a mass scale $M_{l}(z)$ at a given redshift $z$ by considering a sphere with a comoving Lagrangian radius $r_{l}=\pi/k$ Bullock and Boylan-Kolchin (2017):
+$M_{l}(z)=$ $\frac{4\pi}{3}{r_{l}}^{3}\rho_{m}(z)$ (4)
+$=$ $\frac{4\pi}{3}{r_{l}}^{3}\rho_{crit}(z)\Omega_{m}(z)$
+$=$ $\frac{4\pi}{3}{r_{l}}^{3}$
+$\rho_{crit,0}\left[\Omega_{m,0}({1+z})^{3}+\Omega_{r,0}({1+z})^{4}+\Omega_{L,0}\right]$
+$\frac{\Omega_{m,0}}{\left[\Omega_{m,0}({1+z})^{3}+\Omega_{r,0}({1+z})^{4}+\Omega_{L,0}\right]}$
+$=$ $\frac{4\pi}{3}~{}{r_{l}}^{3}~{}\rho_{crit,0}~{}\Omega_{m,0},$ (5)
+where $\rho_{crit}(z)$,$\rho_{m}(z)$ are the comoving critical and matter densities respectively.
+$\Omega_{m}(z)$ is the cosmological density parameter for matter at redshift z.
+$\Omega_{m,0},\Omega_{r,0},\Omega_{L,0}$ are the cosmological density parameters at $z=0$ for matter, radiation, and dark energy respectively.
+We assume a flat universe so the curvature density parameter at $z=0$ is zero, i.e.
+$\Omega_{k,0}=0$.
+As a result, the mass scale is independent of redshift $z$ because both the comoving Lagrangian radius $r_{l}$ and the comoving average matter density $\rho_{m}(z)$ do not depend on redshift.
+We could further name a pivot mass $M_{p}(k_{p})$ for BT model, structures less massive than which could be enhanced while structures more massive should not be affected:
+$M_{p}(k_{p})=$ $5.29\times 10^{12}\left(\frac{\Omega_{m,0}}{0.3}\right)\left(\frac{H_{0}}{70}\right)^{2}\left(\frac{k_{p}}{1~{}{\rm cMpc}^{-1}}\right)^{-3}{\rm M}_{\odot}.$ (6)
 For the BT_deep and BT_soft models, $M_{p}$ would be $\sim 1.1\times 10^{11}~{}{\rm M}_{\odot}$ and $\sim 1.4\times 10^{13}~{}{\rm M}_{\odot}$ respectively, using the cosmological parameters in our paper (see subsubsection III.2.1).
 The host halo of an MW-size galaxy is believed to be around $M_{\rm halo}\sim 10^{12}~{}{\rm M}_{\odot}$.
-Therefore, while both the BT_deep and BT_soft models can enhance the substructures of MW host size halos, only in BT_soft the main halo can be affected.
+Therefore, while both the BT_deep and BT_soft models can enhance the substructures of MW host mass halos, only in BT_soft the main halo can be affected.
 To illustrate the impact of the BT model on the dark matter halo statistics across different redshifts, we apply a public semi-numerical code genmf Reed et al.
@@
 This will help us explore BT PPS’s impact on nearby galaxies and strong lensing observations.
-Figure 3: The cumulative halo mass function generated by genmf, at z=9 (dashed line) and z=0 (solid line), with PL (orange), BT_deep (green), and BT_soft (red).
-The upper panel shows the cumulative halo mass function per unit comoving volume.
-The bottom panel shows the ratios of BT to PL.
-The triangle symbols on the solid lines show the pivot masses $M_{p}$ for BT_deep (green) and BT_soft (red), respectively.
 III Numerical Methods
@@
 III.1 Zoom-in initial conditions (IC)
-We carry out a suite of cosmological zoom-in simulations of MW host-size halos.
+We carry out a suite of cosmological zoom-in simulations of MW host mass halos.
 We increase the number of particles in a selected zoom-in region, which contains high-resolution (HRS) particles.
@@
 For the convenience of comparison, we adopt the MUSIC configuration files of the existing simulation project FIRE (Hopkins et al., 2018; Wetzel et al., 2023).
-We choose the zoom-in initial conditions at the size of MW host halos (i.e., $M_{h}\sim 10^{12}~{}{\rm M}_{\odot}$ in mass), specifically, m12i (Wetzel et al., 2016) and m12f (Garrison-Kimmel et al., 2017).
+We choose the zoom-in initial conditions at the mass of MW host mass halos (i.e., $M_{h}\sim 10^{12}~{}{\rm M}_{\odot}$ in mass), specifically, m12i (Wetzel et al., 2016) and m12f (Garrison-Kimmel et al., 2017).
 We obtain those configuration files from the Flatiron Institute Data Exploration and Comparison Hub  [^3] Publicly available at https://flathub.flatironinstitute.org/fire.
+Several properties of the m12i and m12f host haloes and simulations are listed in Table 2.
+m12f is 30% more massive and more extended than m12i, resulting in also more substructures.
+m12i is more concentrated than m12f as expected from its lower halo mass (Diemer and Kravtsov, 2015).
+m12i has a series of mergers around redshift $z\gtrsim 1$, while m12f is slightly later forming than m12i.
 III.2 Simulation code
-We perform several zoom-in simulations using the state-of-the-art cosmological simulation code SWIFT (Schaller et al., 2018, 2024) [^4] Publicly available at https://swiftsim.com/.
+We perform several zoom-in simulations using the state-of-the-art cosmological simulation code SWIFT (Schaller et al., 2018, 2024)444 publicly available at https://swiftsim.com/.
 For N-body simulations, SWIFT uses the Fast Multipole Method (FMM) at small scales (Dehnen, 2000), which can reduce the complexity of gravity calculation (including both the potential and force calculation) to $O(N)$.
@@
   2. 2.
-    Hubble constant $H_{0}=70.2\ \rm{km}\ \rm{s}^{-1}\ \rm{Mpc}^{-1}$
+    Hubble constant $H_{0}=70.2\ \rm{km}\ \rm{s}^{-1}\ \rm{cMpc}^{-1}$
   3. 3.
@@
   4. 4.
-    Root-mean-square matter fluctuation averaged over a sphere of radius $8h^{-1}~{}\rm{Mpc}$, $\sigma_{8}=0.807$.
+    Root-mean-square matter fluctuation averaged over a sphere of radius $8h^{-1}~{}\rm{cMpc}$, $\sigma_{8}=0.807$.
 III.3 Halo finder and halo property post-processor
 Current halo finders can be categorized into three types based on working mechanism: (A) the configuration space finders, (B) the phase space finders (which utilize both spatial and velocity space information), and (C) the tracking finders (which build the particle list of a certain subhalo based on its progenitors).
-In the main text of the paper, we adopt HBT+HERONS (Han et al., 2017) (Forouhar Moreno et al in preparation)  [^5] This is an updated version of HBT+ that improves the tracking of subhalos.
+In the main text of the paper, we adopt HBT-HERONS (Han et al., 2017; Forouhar Moreno et al., 2025)  [^5] This is an updated version of HBT+ that improves the tracking of subhalos.
 It can be found at https://github.com/SWIFTSIM/HBT-HERONS as our halo finder, which is a tracking finder.
@@
 We choose the position of the most bound particle as the halo center.
-We use SOAP (Spherical Overdensity and Aperture Processor) [^6] https://github.com/SWIFTSIM/SOAP to calculate the spherical overdensity properties, based on the results from HBT+HERONS.
+We use SOAP (Spherical Overdensity and Aperture Processor; McGibbon et al.
+submitted) [^6] https://github.com/SWIFTSIM/SOAP to calculate the spherical overdensity properties for the results from halo finders.
 III.3.1 Convergence radius
@@
     $V_{\rm cir}(R)=\sqrt{\frac{GM(<R)}{R}},$ (8)
-    where $R$ is the distance to the halo center, $M(<R)$ the enclosed mass that is bound to the halo, within radius $R$.
+    where $R$ is the distance to the halo center and $M(<R)$ the enclosed mass that is bound to the halo, within radius $R$.
 The $R-V_{\rm cir}(R)$ graph is called the galaxy rotation curve, and $V_{\rm max}$ is simply the maximum circular velocity of the rotation curve.
-In the simulation, the subhalo would lose its mass due to the tidal forces from the host halo.
-However, $V_{\rm max}$ is robust and representative of a dark matter subhalo because it reflects the subhalo’s central, gravitationally bound part, which is less affected by the stripping effect and how the boundary of the subhalo is defined.
+In the simulation, the subhalo can lose its mass due to tidal forces from the host halo.
+However, $V_{\rm max}$ is a more robust measurement of subhalo mass, since it reflects the subhalo’s central, gravitationally bound part, which is less sensitive to how the boundary of the subhalo is defined.
     During the evolution history of a dark matter halo, its $V_{\rm max}$ may change, and the largest $V_{\rm max}$ is called the peak circular velocity, $V_{\rm peak}$.
@@
 The color scales are the same across the three panels.
-All the images are 2D plane projections of cubic volumes with a side length of $400~{}\rm{kpc}$, centered at the main halo’s most bound particle identified by HBT+HERONS.
+All the images are 2D plane projections of cubic volumes with a side length of $400~{}\rm{kpc}$, centered at the main halo’s most bound particle identified by HBT-HERONS.
 Table 2 shows some selected properties for the main halos in all the simulations.
 We learn that the main halo of BT_soft is larger than that of the PL model in mass, radius, concentration, and $V_{\rm max}$.
-However, the BT_deep model behaves similarly to the PL model, consistent with the pivot mass $M_{p}$ based on Equation 5.
+However, the BT_deep model behaves similarly to the PL model, because the mass of the halo is well above the pivot mass of $M_{p}$ based on Equation 6.
 Figure 4 shows the dark matter projection maps of the PL, BT_deep and BT_soft m12i simulations.
@@
 (2019).
-In the rest of this subsection, we would quantify these differences in terms of the subhalo mass function, the subhalo $V_{\rm max}$ function (velocity refers to $V_{\rm max}$ of the subhalos), the subhalo radial distribution (subhalo distance function), the cumulative substructure mass fraction, the subhalo $R_{\rm max}-V_{\rm max}$ relationship, and the main halo radial density profile.
+In the rest of this subsection, we quantify these differences in terms of the main halo radial density profile, the subhalo mass function, the subhalo $V_{\rm max}$ function, the subhalo radial distribution (subhalo distance function), the cumulative substructure mass fraction and the subhalo $R_{\rm max}-V_{\rm max}$ relationship.
 IV.2.1 Main Halo Radial Density Profile
@@
 We show the NFW fitting result in dashed lines and our simulation results in solid lines.
-The upper panels of Figure 5 show a good match between the simulation results and the NFW fitting lines.
+The upper panels of Figure 5 777 For m12f, the main halo center identified by HBT-HERONS is shifted from the real center by $\leq 2{\rm kpc}$, resulting from a bug existing in the code version we used (see Appendix A7 of Forouhar Moreno et al.
+(2025)), but which has since subsequently been fixed.
+Given this, we adopt the better centering offered by another halo finder, VELOCIraptor, Elahi et al.
+(2019) in this figure.
+Other figures are not affected by this small shift.
+show a good match between the simulation results and the NFW fitting lines.
 Thus we can use the concentration parameters given by the NFW profile to describe the radial density profile.
@@
 Though both m12i and m12f are using the same parameters in the halo finder and thus the same mass of the minimum size halo, m12f has a larger $M_{200c}$ so its scaled subhalo mass limit $\mu=M_{\rm sub}/M_{200c}$ would be smaller.
+Figure 6: The cumulative scaled subhalo mass functions for m12i (colored solid) and m12f (colored dashed) respectively.
+The upper panel shows the subhalo mass function for all the subhalos within $R_{50c}$ from the main halo center in the PL (orange), BT_deep (green), and BT_soft (red) simulations, with the subhalo mass $M_{\rm sub}$ scaled by $M_{200c}$ of the main halo.
+The bottom panel shows the ratios of the BT to PL numbers.
+The shaded areas show the subhalo mass cutoff due to the halo finder’s setting for minimum halo size, where we set it to 20 particles.
+However, based on our resolution study in Appendix B, the subhalo mass function is reliable on $\mu\geq 10^{-5.5}$, i.e.
+for subhalos with at least 50 particles.
+$\mu\sim 10^{-5.5}$ is also approximately the $\mu$ point where the subhalo mass function starts to deviate from a power-law function, i.e.
+a straight line in the log-log graph.
 In the bottom panel, the BT models have an apparent enhancement in the number of dark matter subhalos at different masses.
@@
 $x=4\cdot\log_{10}\frac{\mu}{\mu_{c}}.$ (10)
-Here, $m$ indicates the maximum relative ratio the BT model can touch compared to the PL model, and $\mu_{c}$ is the middle point between $\mu$ starting to drop and $\mu$ value finishing the dropping.
+Figure 7: The cumulative scaled subhalo $V_{\rm max}$ functions for m12i (colored solid) and m12f (colored dashed) respectively.
+The upper panel shows the subhalo $V_{\rm max}$ function for all the subhalos within $R_{100c}$ from the main halo center in the PL (orange), BT_deep (green), and BT_soft (red) simulations, with $V_{\rm max}$ of the subhalo scaled by $V_{200c}$ of the main halo.
+The bottom panel shows the ratios of the BT to PL numbers.
+The shaded areas show the unresolved $V_{\rm max}$ range, the limit of which is the median $V_{\rm max}$ value among all the subhalos with 100 particles in the PL simulations.
+100 is an empirical lower limit of particle number for dark matter halo to be resolved (Rodríguez-Puebla et al., 2016).
+Our resolution limit study in Appendix B also shows that for subhalos with at least 100 particles, the relative deviation in the subhalo $V_{\rm max}$ function does not exceed 30%, which is within acceptable limits.
+Here, $m$ indicates the maximum relative ratio the BT model can reach compared to the PL model, and $\mu_{c}$ is the middle point between $\mu$ starting to drop and $\mu$ value finishing the dropping.
 According to the definition of Equation 9, $\mu_{c}$ is also where the inverse S shape function drops to half.
@@
 After getting this fitting function from the bottom ratio panel, we draw it on the upper panel of Figure 6, showing a good fit with the simulation results.
-Figure 6: The cumulative scaled subhalo mass functions for m12i (colored solid) and m12f (colored dashed) respectively.
-The upper panel shows the subhalo mass function for all the subhalos within $R_{50c}$ from the main halo center in the PL (orange), BT_deep (green), and BT_soft (red) simulations, with the subhalo mass $M_{\rm sub}$ scaled by $M_{200c}$ of the main halo.
-The bottom panel shows the ratios of the BT to PL numbers.
-The shaded areas show the subhalo mass cutoff due to the halo finder’s setting for minimum halo size, where we set it to 20 particles.
-However, based on our resolution study in Appendix B, the subhalo mass function is reliable on $\mu\geq 10^{-5.5}$, i.e.
-for subhalos with at least 50 particles.
-$\mu\sim 10^{-5.5}$ is also approximately the $\mu$ point where the subhalo mass function starts to deviate from a power-law function, i.e.
-a straight line in the log-log graph.
 IV.2.3 Subhalo $V_{\rm max}$ Function
@@
 This indicates that the BT_soft model has a stronger enhancement for a wider range of halo sizes than the BT_deep model.
-Figure 7: The cumulative scaled subhalo $V_{\rm max}$ functions for m12i (colored solid) and m12f (colored dashed) respectively.
-The upper panel shows the subhalo $V_{\rm max}$ function for all the subhalos within $R_{100c}$ from the main halo center in the PL (orange), BT_deep (green), and BT_soft (red) simulations, with $V_{\rm max}$ of the subhalo scaled by $V_{200c}$ of the main halo.
-The bottom panel shows the ratios of the BT to PL numbers.
-The shaded areas show the unresolved $V_{\rm max}$ range, the limit of which is the median $V_{\rm max}$ value among all the subhalos with 100 particles in the PL simulations.
-100 is an empirical lower limit of particle number for dark matter halo to be resolved (Rodríguez-Puebla et al., 2016).
-Our resolution limit study in Appendix B also shows that for subhalos with at least 100 particles, the relative deviation in the subhalo $V_{\rm max}$ function does not exceed 30%, which is within acceptable limits.
 IV.2.4 Subhalo Radial Distribution
@@
 As suggested by Springel et al.
-(2008), within the range of $10^{5}~{}{\rm M}_{\odot}\sim 10^{10}~{}{\rm M}_{\odot}$, massive subhalos tend to exist in the radially outer region while different mass bins follow the same line.
-Both our BT and PL simulation results also show this trend.
-Although the BT subhalo number density is higher, the radial range where subhalos exist within that mass range does not change.
+(2008), the subhalo radial normalized number density in the PL model follows a universal profile across different mass bins, i.e.
+the grey Aquarius A1 line in the four panels of Figure 8.
+However, in more massive bins, the subhalos appear to exist only in the outer region.
+It is likely due to the low number density of subhalos near the center Springel et al.
+(2008) and lack of statistics using only two zoom-in simulations.
 Figure 8: The subhalo radial number density profiles at four different mass bins for m12i (colored solid) and m12f (colored dashed).
@@
 (2018) suggested there could be a “too many satellites” problem in the nearby universe.
-These motivate us to consider alternative cosmological models with an enhanced Primordial Power Spectrum at small scales ($>1~{}{\rm Mpc}^{-1}$), namely the blue-tilted (BT) model.
-In this study, we have conducted cosmological zoom-in simulations of MW host-size halos with the blue-tilted primordial power spectrum.
+These motivate us to consider alternative cosmological models with an enhanced Primordial Power Spectrum at small scales ($>1~{}{\rm cMpc}^{-1}$), namely the blue-tilted (BT) model.
+In this study, we have conducted cosmological zoom-in simulations of MW host mass halos with the blue-tilted primordial power spectrum.
 We have considered two blue-tilted models from Ref.
-Hirano and Yoshida (2024): BT_soft with the enhancement at $\gtrsim 1~{}{\rm Mpc}^{-1}$ and BT_deep with the enhancement at $\gtrsim 4~{}{\rm Mpc}^{-1}$.
+Hirano and Yoshida (2024): BT_soft with the enhancement at $\gtrsim 1~{}{\rm cMpc}^{-1}$ and BT_deep with the enhancement at $\gtrsim 4~{}{\rm cMpc}^{-1}$.
 They are all within the accepted parameter space according to the JWST observation Hirano and Yoshida (2024) and the dwarf galaxy central density observation Dekker and Kravtsov (2024).
@@
 We acknowledge the support of the CUHK Central High-Performance Computing Cluster, on which the computation in this work has been performed; and we thank Edward So, Anny Cheung, and Nicky Leung for their assistance.
-TKC is supported by the ‘Improvement on Competitiveness in Hiring New Faculties’ Funding Scheme from the Chinese University of Hong Kong (4937210, 4937211, 4937212).
+TKC is supported by the ‘Improvement on Competitiveness in Hiring New Faculties’ Funding Scheme from the Chinese University of Hong Kong (4937210, 4937211, 4937212) and RGC Early Career Scheme (24301524).
 TKC and JW are partially supported by the Direct Grant project (4053662) from the Chinese University of Hong Kong.
@@
 Appendix A Results from another halo finder: VELOCIraptor
-Apart from HBT+HERONS, we also identify the halos and subhalos with VELOCIraptor (VR) (Elahi et al., 2019), a code based on the friends-of-friends (FOF) algorithm that identifies peaks in 6D phase-space to find subhalos (both the configuration space and phase space).
-We compare the HBT+HERONS and VR results on the subhalo mass function, subhalo $V_{\rm max}$ function, and subhalo radial distribution in subsection A.1.
+Apart from HBT-HERONS, we also identify the halos and subhalos with VELOCIraptor (VR) (Elahi et al., 2019), a code based on the friends-of-friends (FOF) algorithm that identifies peaks in 6D phase-space to find subhalos (both the configuration space and phase space).
+We compare the HBT-HERONS and VR results on the subhalo mass function, subhalo $V_{\rm max}$ function, and subhalo radial distribution in subsection A.1.
 Currently, SOAP (see subsection III.3) only allows the calculation of spherical overdensity properties for field halos.
-Besides, HBT+HERONS (without SOAP post process) could only calculate limited kinds of SO properties.
+Besides, HBT-HERONS (without SOAP post process) could only calculate limited kinds of SO properties.
 Hence, we used VR (without SOAP post process) to calculate the virial masses for all the halos (also including subhalos) and show the $M_{\rm vir}-V_{\rm max}$ relationship in subsection A.2.
@@
 Otherwise, the subhalos near the center of the main halo will have unphysically large masses.
-Our essential parameters for VR are listed in footnote 11, based on the sample configuration for 6D dark matter-only simulation  [^9] https://velociraptor-stf.readthedocs.io/en/latest/_downloads/97ba16973bcff9d4afe940c2e13d12d2/sample_dmcosmological_6dfof_subhalo.cfg.
-We set the most bound particle’s position as the reference frame center to be consistent with HBT+HERONS for comparison in subsection A.1.
+Our essential parameters for VR are listed in footnote 12, based on the sample configuration for 6D dark matter-only simulation  [^10] https://velociraptor-stf.readthedocs.io/en/latest/_downloads/97ba16973bcff9d4afe940c2e13d12d2/sample_dmcosmological_6dfof_subhalo.cfg.
+We set the most bound particle’s position as the reference frame center to be consistent with HBT-HERONS for comparison in subsection A.1.
 Parameters Value
@@
 For detailed explanations of these parameters, refer to the usage section in the documentation  [^12] https://velociraptor-stf.readthedocs.io/en/latest/usage.html.
-A.1 Comparison between halo finders: HBT+HERONS and VELOCIraptor
+A.1 Comparison between halo finders: HBT-HERONS and VELOCIraptor
 Here, we choose m12i to analyze the difference between the two halo finders.
 VR stands for VELOCIraptor in both the text and figures.
-In this subsection, we also use SOAP to process VR and HBT+HERONS catalogs in a consistent manner.
+In this subsection, we also use SOAP to process VR and HBT-HERONS catalogs in a consistent manner.
 However, for the previous literatures we fit with, COCO (Hellwing et al., 2016) and Aquarius (Springel et al., 2008) use SUBFIND (Springel et al., 2001) as their halo finder, which is a configuration space finder; Cautun fit line (Cautun et al.
@@
 The relative ratio between the two halo finders would be more meaningful.
-Figure 11: The cumulative scaled subhalo mass functions for m12i, with two halo finders HBT+HERONS (colored solid) and VR (colored dashed), respectively.
+Figure 11: The cumulative scaled subhalo mass functions for m12i, with two halo finders HBT-HERONS (colored solid) and VR (colored dashed), respectively.
 The upper panel shows the subhalo mass function for the PL (orange), BT_deep (green), and BT_soft (red) models, with the subhalo mass $M_{\rm sub}$ scaled by $M_{200c}$ of the main halo.
-The bottom panel shows the ratios of the VR to HBT+HERONS number.
+The bottom panel shows the ratios of the VR to HBT-HERONS number.
 The shaded areas are the same as that in Figure 6.
 The shaded areas overlap because we set the minimum size of the subhalo to be the same in both halo finders.
-The subhalo mass function comparison between VR and HBT+HERONS is shown in Figure 11.
+The subhalo mass function comparison between VR and HBT-HERONS is shown in Figure 11.
 These two halo finders agree with each other for $10^{-5}<\mu<10^{-3},10^{7}~{}{\rm M}_{\odot}<M_{\rm sub}<10^{9}~{}{\rm M}_{\odot}$.
-However, HBT+HERONS is higher and closer to the COCO fitting line than VR at the high mass end, implying that HBT+HERONS is better at finding more massive subhalos.
-But at the low mass end, VR appears to have more subhalos than HBT+HERONS.
-Figure 12: The cumulative scaled subhalo $V_{\rm max}$ functions for m12i, with two halo finders HBT+HERONS (colored solid) and VR (colored dashed), respectively.
+However, HBT-HERONS is higher and closer to the COCO fitting line than VR at the high mass end, implying that HBT-HERONS is better at finding more massive subhalos.
+But at the low mass end, VR appears to have more subhalos than HBT-HERONS.
+Figure 12: The cumulative scaled subhalo $V_{\rm max}$ functions for m12i, with two halo finders HBT-HERONS (colored solid) and VR (colored dashed), respectively.
 The upper panel shows the subhalo $V_{\rm max}$ function for the PL (orange), BT_deep (green), and BT_soft (red) models, with the subhalo $V_{\rm max}$ scaled by $V_{200c}$ of the main halo.
-The bottom panel shows the ratios of the VR to HBT+HERONS number.
+The bottom panel shows the ratios of the VR to HBT-HERONS number.
 The shaded areas are the same as those in Figure 7.
-The subhalo $V_{\rm max}$ function comparison between VR and HBT+HERONS is shown in Figure 12.
+The subhalo $V_{\rm max}$ function comparison between VR and HBT-HERONS is shown in Figure 12.
 These two halo finders give similar results except at the low $V_{\rm max}$ end near the resolution limit.
-When approaching the limit $V_{\rm max}$ value, VR deviates from the power-law behavior later than HBT+HERONS.
-Figure 13: The subhalo radial number density profiles at four different mass bins for m12i, using different halo finders HBT+HERONS (colored solid) and VR (colored dashed).
-We consider the PL (orange), BT_deep (green), and BT_soft (red) models in each subfigure (also in each mass bin): the upper panel is the same definition as that in Figure 8; the bottom panel shows the ratios of the VR results to the HBT+HERONS results.
-The subhalo radial distribution comparison between VR and HBT+HERONS is shown in Figure 13.
-In mass bins larger than $10^{7}~{}{\rm M}_{\odot}$, VR and HBT+HERONS have nearly the same results.
-But, at the lower mass end, VR has much fewer subhalos in the inner region than HBT+HERONS and the Aquarius result.
-To summarize, VR and HBT+HERONS behave similarly for the subhalo mass and $V_{\rm max}$ functions.
+When approaching the limit $V_{\rm max}$ value, VR deviates from the power-law behavior later than HBT-HERONS.
+Figure 13: The subhalo radial number density profiles at four different mass bins for m12i, using different halo finders HBT-HERONS (colored solid) and VR (colored dashed).
+We consider the PL (orange), BT_deep (green), and BT_soft (red) models in each subfigure (also in each mass bin): the upper panel is the same definition as that in Figure 8; the bottom panel shows the ratios of the VR results to the HBT-HERONS results.
+The subhalo radial distribution comparison between VR and HBT-HERONS is shown in Figure 13.
+In mass bins larger than $10^{7}~{}{\rm M}_{\odot}$, VR and HBT-HERONS have nearly the same results.
+But, at the lower mass end, VR has much fewer subhalos in the inner region than HBT-HERONS and the Aquarius result.
+To summarize, VR and HBT-HERONS behave similarly for the subhalo mass and $V_{\rm max}$ functions.
 However, for the radial properties (the subhalo radial distribution function), VR gaves a lower halo count near the host halo center for the low halo mass bin.